### v3

# references.bib

Abstract.
In this paper, we prove the integrality conjecture for quotient stacks given by a symmetric representation of a reductive group. This result will be crucial to generalise cohomological Donaldson–Thomas theory of $2$ and $3$-Calabi–Yau categories to 00 and $(-1)$-shifted symplectic stacks respectively.
1. Introduction
Given a connected reductive group $G$ and a representation $V$ of $G$, the quotient stack $\mathfrak{X}=V/G$ and the GIT quotient $V/\!\!/G$ are objects of interest in geometry and representation theory. In the context of quiver representations, the geometry and topology of the GIT quotient $V/\!\!/G$ has been studied deeply by Reineke, see [reineke2008moduli] for a survey. A particularly nice case is when $V$ is a symmetric representation of $G$, in the sense that the weights of $V$ counted with multiplicities are invariant under the involution $\alpha\mapsto-\alpha$.
In [vspenko2017non, vspenko2021semi], the authors study semiorthogonal decompositions of the coherent derived category $\mathcal{D}(V/G)$ of $V/G$ in order to obtain noncommutative resolutions of $V/\!\!/G$. The tools are the window categories defined and studied by Halpern-Leistner [halpern2015derived]. Their formalism deals with quasi-symmetric representations $V$ of a reductive group $G$. Quasi-symmetricity is the vanishing property of the sum of weights of $V$ on any given line through the origin of the character lattice and is weaker than the symmetricity.
In this paper, we are interested in the interplay between the topology of $V/G$ and that of $V/\!\!/G$. More precisely, we give a decomposition of the (infinite dimensional) cohomology vector space $\mathrm{H}^{*}(V/G)$ of the quotient stack $V/G$ in terms of a finite number of finite-dimensional cohomologically graded vector spaces $\mathcal{P}_{\lambda}$ indexed by cocharacters $\lambda\in X_{*}(T)$ of a maximal torus $T$ of $G$, using parabolic induction (Theorem 1.1), when $V$ is symmetric. Such a result is often called cohomological integrality. The refined Donaldson–Thomas invariant of $V/G$ are then defined as the Betti numbers of the vector spaces $\mathcal{P}_{\lambda}$. In this paper, we initiate the study of these enumerative invariants, which encode representation-theoretic information about $V/G$.
In geometry, stacks of the form $V/G$ and their good moduli space $V/\!\!/G$ give, by Luna étale slice theorems [alper2020luna], local models for smooth stacks $\mathfrak{X}$ with a good moduli space $\mathcal{X}$. Therefore, this work is the ground for the study of the geometry and topology of morphisms $\mathfrak{X}\rightarrow\mathcal{X}$ from a symmetric smooth stack to its good moduli space.
When the group $G$ is a product of general linear groups and $V$ is a symmetric representation of $G$ given by the representation space of a symmetric quiver $Q$ (a quiver having as many arrows in both possible directions between any two vertices), the cohomological integrality isomorphism has been proven by Efimov in [efimov2012cohomological], answering a conjecture of Kontsevich and Soibelman [kontsevich2011cohomological, Conjecture 1]. It is formulated in terms of the cohomological Hall algebra of $Q$ [kontsevich2011cohomological]. The cohomological integrality of Efimov is the building block of all cohomological integrality results for categories of homological dimension one, and for $2$ and $3$-Calabi–Yau categories [meinhardt2019donaldson, davison2020cohomological, davison2022bps, davison2023bps]. These integrality theorems are one of the most sought-after theorems in the theory of cohomological Hall algebras, as they provide strong structural results and deep connections between the topology of the stack and the moduli space. The integrality also gives cohomologically refined enumerative invariants whose study is of algebro-geometric interest.
In particular, the cohomological integrality isomorphisms for the $2$-Calabi–Yau categories of semistable Higgs bundles on a smooth projective curve and twisted local systems is at the heart of nonabelian Hodge isomorphisms for stacks [davison2023nonabelian, davison2022bps, hennecart2023nonabelian]. Nonabelian Hodge isomorphisms for stacks are crucial to study the $\mathrm{P}=\mathrm{W}$ conjecture for $\mathrm{GL}_{n}$ for the singular moduli spaces.
A $2$-Calabi–Yau category $\mathscr{C}$ gives rise to a 00-shifted symplectic stack while a $3$-Calabi–Yau category $\mathscr{C}$ provides us with a $(-1)$-shifted symplectic stack [pantev2013shifted]. In the $2$-Calabi–Yau case, fully settled in [davison2022bps, davison2023bps], the cohomological integrality gives a decomposition of the Borel Moore homology $\mathrm{H}^{\mathrm{BM}}_{*}(\mathfrak{M}_{\mathscr{C}})$ of (a substack of) the stack of objects in the category $\mathscr{C}$ in terms of the intersection cohomology $\mathrm{IH}(\mathcal{M}_{\mathscr{C}})$ of the good moduli space $\mathcal{M}_{\mathscr{C}}$ of $\mathfrak{M}_{\mathscr{C}}$. The study of cohomological integrality for $2$-Calabi–Yau categories has several applications, such as the study of the cohomology of Nakajima quiver varieties and cuspidal polynomials of quivers [davison2023bps] and the comparison between cohomological Hall algebras and Maulik–Okounkov Yangians [maulik2019quantum, botta2023okounkov, schiffmann2023cohomological].
In the $3$-Calabi–Yau case, the cohomological integrality is, in full generality, a conjecture recalled in [davison2023bps, Conjecture 1.9]. It is proven for quivers with potential in [davison2020cohomological]. It concerns the study of the critical cohomology of the moduli stack of objects in the category. The critical cohomology is the cohomology of a perverse sheaf, which is either the sheaf of vanishing cycles or a globalisation of it [ben2015darboux, brav2015symmetries, brav2019darboux, kiem2012categorification]. In the case of general $3$-Calabi–Yau categories, significant progress towards the integrality conjecture is made in [kinjo2024cohomological].
All known cohomological integrality isomorphisms are expressed in terms of cohomological Hall algebras, which is a convenient and powerful way to gather the data of parabolic inductions. If rather than a $2$-Calabi–Yau category, one considers a 00-shifted symplectic stack $\mathfrak{M}$, one expects to have similar cohomological integrality isomorphisms. However, the cohomological Hall algebra structure is not available and one has to deal directly with the parabolic induction morphisms. This is what we do in this paper in the situation of smooth symmetric stacks given by a symmetric representation of a reductive group. In particular, the very definition of the cohomological integrality morphism (see Theorem 1.1) had to be given.
There are motivic versions of integrality results, studied for example in [bu2024motivic] following [kontsevich2008stability] and motivic Donaldson–Thomas theory [kontsevich2010motivic, meinhardt2017introduction, bussi2019motivic].
1.1. Main results
In this paper, we work over an algebraically closed field of characteristic zero.
1.1.1. Induction
Let $G$ be a connected reductive group and $V$ a representation of $G$. We let $T\subseteq G$ be a maximal torus. For any cocharacter $\lambda\colon\mathbf{G}_{\mathrm{m}}\rightarrow T$, we let $G^{\lambda}\subseteq G$ be the corresponding Levi subgroup of $G$ [digne2020representations, Proposition 1.22] and $V^{\lambda}$ the representation of $G^{\lambda}$ given by the $\lambda$-fixed subspace of $V$. If $\lambda=0$ is the trivial cocharacter, we write $V=V^{0}$ and $G=G^{0}$.
There is an induction map
(1.1) $\mathrm{Ind}_{\lambda}\colon\mathrm{H}^{*+d_{\lambda}}(V^{\lambda}/G^{\lambda})\rightarrow\mathrm{H}^{*+d_{0}}(V/G),$
where $d_{\lambda}=\dim V^{\lambda}-\dim G^{\lambda}$2.2). If $V$ is symmetric (i.e. $V$ and $V^{*}$ have the same sets of weights, Definition 2.3), then the induction map preserves cohomological degrees (Lemma 2.5).
We define an order on the set $X_{*}(T)$ of cocharacters of $T$ as follows:
$\lambda\preceq\mu\stackrel{{\scriptstyle\text{def.}}}{{\iff}}\left\{\begin{aligned} &V^{\lambda}\subseteq V^{\mu}\\ &\mathfrak{g}^{\lambda}\subseteq\mathfrak{g}^{\mu}.\end{aligned}\right.$
We let $\mathscr{P}_{V}\coloneqq X_{*}(T)/\sim$ where $\sim$ is the equivalence relation given by $\lambda\sim\mu\iff(\lambda\preceq\mu\text{ and }\mu\preceq\lambda)$. In other words, two cocharacters are equivalent if their fixed-point sets inside $V$ and $\mathfrak{g}$ both coincide. The set $\mathscr{P}_{V}$ is finite. If $V$ is symmetric, the induction map $\mathrm{Ind}_{\lambda}$ only depends on the class $\overline{\lambda}\in\mathscr{P}_{V}$ up to a sign (Lemma 2.14).
We let $W=N_{G}(T)/T$ be the Weyl group of $G$. The natural action of $W$ on $X_{*}(T)$ descends to a $W$-action on $\mathscr{P}_{V}$2.2). The image of the induction map $\mathrm{Ind}_{\lambda}$ only depends on the class $\widetilde{\lambda}\in\mathscr{P}_{V}/W$ of $\lambda$ (Lemma 2.15).
1.1.2. Cohomological integrality
Let $G$ be a reductive group with maximal torus $T$ and $V$ be a finite dimensional symmetric representation of $G$ (Definition 2.3). For any $\overline{\lambda}\in\mathscr{P}_{V}$, we let $\lambda\in X_{*}(T)$ be an arbitrary lift. For $\widetilde{\lambda}\in\mathscr{P}_{V}/W$, we let $\lambda\in X_{*}(T)$ be an arbitrary lift.
For $\lambda\in X_{*}(T)$, we let $W_{\lambda}=\{w\in W\mid V^{w\cdot\lambda}=V^{\lambda}\text{ and }\mathfrak{g}^{w\cdot\lambda}=\mathfrak{g}^{\lambda}\}$. This is a subgroup of $W$ (Lemma 4.1).
We let $G_{\lambda}\subseteq G$ be the intersection of the kernel of the action map $G^{\lambda}\rightarrow\mathrm{GL}(V^{\lambda})$ with the center of $G^{\lambda}$ and $\mathfrak{g}_{\lambda}\coloneqq\mathrm{Lie}(G_{\lambda})$. Note that the group $G_{\lambda}$ is not necessarily connected, but this can essentially be ignored (Lemma 4.19).
For $\alpha\in X^{*}(T)$, we let $V_{\alpha}\coloneqq\{v\in V\mid\forall t\in T,t\cdot v=\alpha(t)\}$ be the $\alpha$-weight space of $V$. We define $\mathfrak{g}_{\alpha}$ similarly.
We let
$k_{\lambda}\coloneqq\frac{\prod_{\alpha\in X^{*}(T),\langle\lambda,\alpha\rangle>0}\alpha^{\dim V_{\alpha}}}{\prod_{\alpha\in X^{*}(T),\langle\lambda,\alpha\rangle>0}\alpha^{\dim\mathfrak{g}_{\alpha}}}.$
We may call $k_{\lambda}$ the induction kernel. We define $\varepsilon_{V,\lambda}\colon W_{\lambda}\rightarrow\{\pm 1\}$ to be the sign such that for any $w\in W_{\lambda}$, $w(k_{\lambda})=\varepsilon_{V,\lambda}(w)k_{\lambda}$ (Proposition 4.8).
Theorem 1.1 (Proposition 4.18 + Corollary 4.27).
There exists cohomologically graded vector spaces $\mathcal{P}_{\lambda}\subseteq\mathrm{H}^{*+d_{\lambda}}(V^{\lambda}/G^{\lambda})$ with a $W_{\lambda}$-action, $\lambda\in X_{*}(T)$, such that the map
$\bigoplus_{\widetilde{\lambda}\in\mathscr{P}_{V}/W}(\mathcal{P}_{\lambda}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda}))^{\varepsilon_{V,\lambda}}\xrightarrow{\sum_{\widetilde{\lambda}\in\mathscr{P}_{V}/W}\mathrm{Ind}_{\lambda}}\mathrm{H}^{*}(V/G)$
induced by the induction morphisms (1.1), is a graded isomorphism, where $\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})\cong\mathrm{Sym}(\mathfrak{g}_{\lambda}^{*})$, $\mathfrak{g}_{\lambda}^{*}$ sits in degree $2$, $\varepsilon_{V,\lambda}\colon W_{\lambda}\rightarrow\{\pm 1\}$ is a character and $(\mathcal{P}_{\lambda}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda}))^{\varepsilon_{V,\lambda}}$ denotes the $\varepsilon_{V,\lambda}$-isotypic component for the natural $W_{\lambda}$-action. Moreover, the vector spaces $\mathcal{P}_{\lambda}$ are finite-dimensional and the inclusion $\mathcal{P}_{\lambda}\subseteq\mathrm{H}^{*+d_{\lambda}}(V^{\lambda}/G^{\lambda})$ factors through the inclusion $\mathrm{H}^{*+d_{\lambda}}(V^{\lambda}/(G^{\lambda}/G_{\lambda}))\rightarrow\mathrm{H}^{*+d_{\lambda}}(V^{\lambda}/G^{\lambda})$.
This type of theorems is sometimes referred to as cohomological integrality. We can reconstruct the infinite dimensional vector space $\mathrm{H}^{*}(V/G)$ using a finite number of finite-dimensional vector spaces $\mathcal{P}_{\lambda}$, $\widetilde{\lambda}\in\mathscr{P}_{V}/W$.
Remark 1.2.
Let $Q=(Q_{0},Q_{1})$ be a symmetric quiver with set of vertices $Q_{0}$ and set of arrows $Q_{1}$. Let $\mathbf{d}\in\mathbf{N}^{Q_{0}}$ be a dimension vector. If $V_{\mathbf{d}}$ is the representation space of $\mathbf{d}$-dimensional representations of $Q$ which is acted on by $\mathrm{GL}_{\mathbf{d}}\coloneqq\prod_{i\in Q_{0}}\mathrm{GL}_{\mathbf{d}_{i}}$, then the isomorphism given by Theorem 1.1 essentially recovers [efimov2012cohomological, Theorem 1.1]. In loc.cit, it was convenient for the author to twist the cohomological Hall algebra multiplication to make it supercommutative, as in [kontsevich2011cohomological]. This twist does not change the images of the induction/multiplication morphisms which means that it is possible to have cohomological integrality isomorphisms even without this twist, although we cannot identify the untwisted CoHA with a supercommutative algebra anymore. We can only identify the underlying vector spaces.
Definition 1.3.
We define the cohomologically refined Donaldson–Thomas invariants of $V/G$ as the dimensions $p_{\lambda,i}\coloneqq\dim\mathcal{P}_{\lambda}^{i}$ of the cohomological degree $i$ piece of the vector space $\mathcal{P}_{\lambda}$ for $\lambda\in X_{*}(T)$. We may call the Euler characteristics $p_{\lambda}\coloneqq\sum_{i\in\mathbf{Z}}(-1)^{i}p_{\lambda,i}$ of $\mathcal{P}_{\lambda}$ the Donaldson–Thomas invariants of $V/G$.
1.1.3. Vanishing bound
The vector spaces $P_{\lambda}$ are finite dimensional, and we can give a rough bound on the cohomological degrees for which they possibly do not vanish. We let $\mathcal{H}_{\lambda}\coloneqq\mathrm{H}^{*+d_{\lambda}}(V^{\lambda}/G^{\lambda})$, where $d_{\lambda}=\dim V^{\lambda}-\dim G^{\lambda}$.
Proposition 1.4.
The cohomologically graded vector space $P_{\lambda}$ given by Theorem 1.1 are concentrated in cohomological degrees $[\dim G^{\lambda}-\dim V^{\lambda},\dim G^{\lambda}]$, that is $\mathcal{P}_{\lambda}^{k}=0$ unless $k\in[\dim G^{\lambda}-\dim V^{\lambda},\dim G^{\lambda}]$.
1.2. Notations and conventions
  1. $\bullet$
    If $V$ is a representation of a finite group $W$ and $\chi$ an irreducible character of $W$, we let $V^{\chi}$ be the $\chi$-isotypic component of $V$.
  2. $\bullet$
    If $X$ is a complex algebraic variety acted upon by an algebraic group $G$, we let $X/G$ be the quotient stack.
  3. $\bullet$
    The multiplicative group is denoted by $\mathbf{G}_{\mathrm{m}}$. The Abelian group of characters of an algebraic torus $T$ is $\mathrm{Hom}_{\mathbf{Z}}(T,\mathbf{G}_{\mathrm{m}})$. The set of cocharacters of $T$ is the Abelian group $\mathrm{Hom}_{\mathbf{Z}}(\mathbf{G}_{\mathrm{m}},T)$.
  4. $\bullet$
    Let $H\subseteq G$ be algebraic groups and $X$ an $H$-variety. We let $X\times^{H}G\coloneqq(X\times G)/H$ where $H$ acts on $X\times G$ by $h\cdot(x,g)=(h\cdot x,gh^{-1})$. We formula $g\cdot(x,g^{\prime})\coloneqq(x,gg^{\prime})$ gives a $G$-action on $X\times^{H}G$.
  5. $\bullet$
    If $R$ is a domain (e.g. $R=\mathrm{H}^{*}_{T}(\mathrm{pt})$ for some algebraic torus $T$), we denote by $\mathrm{Frac}(R)$ its fraction field.
  6. $\bullet$
    If $G$ is a possibly disconnected algebraic group, then we denote by $G^{0}$ its neutral component.
1.3. Acknowledgements
At various stages of the preparation of this work, the author was supported by the Royal Society and by the National Science Foundation under Grant No. DMS-1928930 and by the Alfred P. Sloan Foundation under grant G-2021-16778, while the author was in residence at the Simons Laufer Mathematical Sciences Institute (formerly MSRI) in Berkeley, California, during the Spring 2024 semester. The author thanks the SLMath Sciences Institute and the University of Edinburgh for the excellent working conditions. The author is grateful to Ben Davison for the support provided, in particular via postdoctoral fellowships, and to Woonam Lim for pointing out a mistake in the example section.
2. Parabolic induction for representations of reductive groups
2.1. The induction diagram
Let $G$ be a reductive group and $V$ a finite-dimensional representation of $G$. For a cocharacter $\lambda\colon\mathbf{G}_{\mathrm{m}}\rightarrow G$, we let $G^{\lambda}\coloneqq\{g\in G\colon\lambda(t)g\lambda(t)^{-1}=g\}$ be the centraliser of $\lambda$, a Levi subgroup of $G$ and $V^{\lambda}\coloneqq\{v\in V\mid\lambda(t)\cdot v=v\}$ be the fixed locus, a representation of $G^{\lambda}$. We also let $P_{\lambda}\coloneqq\{g\in G\mid\lim_{t\rightarrow 0}\limits\lambda(t)g\lambda(t)^{-1}\text{ exists}\}$, a parabolic subgroup of $G$ and $V^{\lambda\geqslant 0}\coloneqq\{v\in V\mid\lim_{t\rightarrow 0}\limits\lambda(t)\cdot v\text{ exists}\}$, a representation of $P_{\lambda}$.
A weight of $V$ is a character $\alpha\colon T\rightarrow\mathbf{G}_{\mathrm{m}}$ such that there exists $v\in V\smallsetminus\{0\}$ such that for any $t\in T$, $t\cdot v=\alpha(t)v$.
We let $\mathcal{W}^{\prime}(V)$ be the set of weights of the representation $V$. We have a direct sum decomposition
$V\cong\bigoplus_{\alpha\in\mathcal{W}^{\prime}(V)}V_{\alpha}$
where $V_{\alpha}$ consists of 00 and nonzero vectors of weight $\alpha$.
We let $\mathcal{W}(V)$ be the collection of weights of $V$, counted multiplicities: $\alpha\in\mathcal{W}^{\prime}(V)$ appears $\dim V_{\alpha}$ times in $\mathcal{W}(V)$.
In particular, we have $\mathcal{W}^{\prime}(\mathfrak{g})$ and $\mathcal{W}(\mathfrak{g})$ where $\mathfrak{g}$ is seen as the adjoint representation of $G$. We have the natural pairing
$\langle-,-\rangle\colon X_{*}(T)\times X^{*}(T)\rightarrow\mathbf{Z}$
between characters and cocharacters. We let $\mathcal{W}^{\lambda>0}(V)\coloneqq\{\alpha\in\mathcal{W}(V)\mid\langle\lambda,\alpha\rangle>0\}$ and we define similarly $\mathcal{W}^{\lambda=0}(V)$, $\mathcal{W}^{\lambda\geqslant 0}(V)$, $\mathcal{W}^{\lambda<0}(V)$ and $\mathcal{W}^{\lambda\leqslant 0}(V)$ and their versions, by forgetting the multiplicities.
It is immediate that $V^{\lambda\geqslant 0}=\bigoplus_{\alpha\in\mathcal{W}^{{}^{\prime}\lambda\geqslant 0}(V)}V_{\alpha}$.
We consider the commutative induction diagram
[…]
where $V/\!\!/G\coloneqq\mathrm{Spec}(\mathbf{C}[V]^{G})$ is the affine GIT quotient of $V$ by $G$.
Lemma 2.1.
The map $q_{\lambda}$ is a vector bundle stack of rank $r_{\lambda}\coloneqq\#\mathcal{W}^{\lambda>0}(V)-\#\mathcal{W}^{\lambda>0}(\mathfrak{g})$ and the map $p_{\lambda}$ is a proper and representable morphism of stacks.
Proof.
The map $p_{\lambda}$ can be presented as the map of quotient stacks $(V^{\lambda\geqslant 0}\times^{P_{\lambda}}G)/G\rightarrow V/G$ coming from the $G$-equivariant map
$\widetilde{p}_{\lambda}\colon\widetilde{V}^{\lambda\geqslant 0}\coloneqq V^{\lambda\geqslant 0}\times^{P_{\lambda}}G\rightarrow V,$
which is projective. Therefore, $p_{\lambda}$ is representable and projective.
The fact that $q_{\lambda}$ is smooth, surjective of relative dimension #$\#\mathcal{W}^{\lambda>0}(V)-\mathcal{W}^{\lambda>0}(\mathfrak{g})$ is easily seen. That it is in addition a vector bundle stack is a rather classical fact (which is not used in this paper). It comes from the facts that $V^{\geqslant\lambda}\rightarrow V^{\lambda}$ is a vector bundle (by the Białinicky-Birula decomposition [bialynicki1973some]) and the map $P_{\lambda}\rightarrow G^{\lambda}$ has kernel the unipotent radical of $P_{\lambda}$. ∎
Lemma 2.2.
The map $\imath_{\lambda}$ is a finite map.
Proof.
The arguments are parallel to that of the proof [meinhardt2019donaldson, Lemma 2.1]. Namely, we consider the commutative diagram
[…]
where $\widetilde{V}^{\lambda\geqslant 0}\coloneqq V^{\lambda\geqslant 0}\times^{P_{\lambda}}G$, so that we have an equivalence of stacks $V^{\lambda\geqslant 0}/P_{\lambda}\simeq\widetilde{V}^{\lambda\geqslant 0}/G$. We now explain the definitions of the maps in this diagram.
The map $q_{\lambda}$ is given by the equivariant map $(V^{\lambda\geqslant 0},P_{\lambda})\rightarrow(V^{\lambda},G^{\lambda})$
Therefore, the pullback of functions by $q_{\lambda}$ induces a morphism
$\overline{q}_{\lambda}^{*}\colon\mathbf{C}[V^{\lambda}]^{G^{\lambda}}\rightarrow\mathbf{C}[\widetilde{V}^{\lambda\geqslant 0}]^{G}$
whose dual is the morphism between affine schemes
$\overline{q}_{\lambda}\colon\mathrm{Spec}(\mathbf{C}[\widetilde{V}^{\lambda\geqslant 0}]^{G^{\lambda}})\rightarrow\mathrm{Spec}(\mathbf{C}[V^{\lambda}])=V^{\lambda}/\!\!/G^{\lambda}.$
The map $\jmath_{\lambda}$ is given by the equivariant morphism $(V^{\lambda},G^{\lambda})\rightarrow(V^{\lambda\geqslant 0},P_{\lambda})$ which induces the morphism
$\overline{\jmath}_{\lambda}^{*}\colon\mathbf{C}[\widetilde{V}^{\lambda\geqslant 0}]^{G}\rightarrow\mathbf{C}[V^{\lambda}]^{G^{\lambda}}$
and dually the morphism of schemes
$\overline{\jmath}_{\lambda}\colon\mathrm{Spec}(\mathbf{C}[V^{\lambda}]^{G^{\lambda}})\rightarrow\mathrm{Spec}(\mathbf{C}[\widetilde{V}^{\lambda\geqslant 0}]^{G}).$
Since $q_{\lambda}\circ\jmath_{\lambda}=\mathrm{id}$, we have $\overline{q}_{\lambda}\circ\overline{\jmath}_{\lambda}=\mathrm{id}$. On the ring of functions, we have $\overline{\jmath}_{\lambda}^{*}\circ\overline{q}_{\lambda}^{*}=\mathrm{id}$ and so $\jmath_{\lambda}^{*}$ is surjective. This tells us that $\overline{\jmath}_{\lambda}$ is a closed immersion. Moreover, $\imath_{\lambda}=\overline{p}_{\lambda}\circ\overline{\jmath}_{\lambda}$ and so it suffices to prove that $\overline{p}_{\lambda}$ is a finite map. Since it is a morphism of finite type complex schemes, it suffices to prove that it is an integral morphism [stacks-project, Lemma 01WJ].
Since $\widetilde{p}_{\lambda}\colon\widetilde{V}^{\lambda\geqslant 0}\rightarrow V$ is a projective morphism (Proof of Lemma 2.1), $(\widetilde{p}_{\lambda})_{*}\mathcal{O}_{\widetilde{V}^{\lambda\geqslant 0}}$ is a coherent $\mathcal{O}_{V}$-module. Therefore, the map $\mathbf{C}[V]\rightarrow\mathbf{C}[\widetilde{V}^{\lambda\geqslant 0}]$ is finite and hence integral [stacks-project, Lemma 01WJ]. We deduce that the map
$\overline{p}_{\lambda}^{*}\colon\mathbf{C}[V]^{G}\rightarrow\mathbf{C}[\widetilde{V}^{\lambda\geqslant 0}]^{G}$
is integral, and hence finite, as follows. Let $a\in\mathbf{C}[\widetilde{V}^{\geqslant\lambda}]^{G}$. It is solution of a monic polynomial with coefficients in $\mathbf{C}[V]$. By applying the Reynolds operator to the equation, we obtain a monic equation with coefficients in $\mathbf{C}[V]^{G}$ of which $a$ is a zero. This concludes.
For $\lambda\in X_{*}(T)$, we define $d_{\lambda}\coloneqq\#\mathcal{W}^{\lambda=0}(V)-\#\mathcal{W}^{\lambda=0}(\mathfrak{g})=\dim V^{\lambda}/G^{\lambda}$. The quantity $r_{\lambda}$ is defined in Lemma 2.1.
Definition 2.3.
We say that a representation $V$ of a reductive group $G$ is symmetric if $V$ and its dual $V^{*}$ have the same sets of weights counted with multiplicities: $\mathcal{W}(V)=\mathcal{W}(V^{*})$.
For later use, we give the following easy lemma.
Lemma 2.4.
Let $V$ be a symmetric representation of $G$. Then,
$d_{\lambda}+2r_{\lambda}=d.$
Proof.
We have
$\mathcal{W}(V)=\mathcal{W}^{\lambda>0}(V)\sqcup\mathcal{W}^{\lambda=0}(V)\sqcup\mathcal{W}^{\lambda<0}(V)$
and by symmetry, #$\#\mathcal{W}^{\lambda>0}(V)=\#\mathcal{W}^{\lambda<0}(V)$. Therefore, #$\#\mathcal{W}(V)=2\#\mathcal{W}^{\lambda>0}(V)+\#\mathcal{W}^{\lambda=0}(V)$. These equalities are also valid for $V=\mathfrak{g}$ under the adjoint action of $G$. We obtain the lemma by combining them together. ∎
2.2. Parabolic induction
For $\lambda\in X_{*}(T)$, we let $\mathbf{Q}_{V^{\lambda}/G^{\lambda}}^{\mathrm{vir}}\coloneqq\mathbf{Q}_{V^{\lambda}/G^{\lambda}}\otimes\mathscr{L}^{-d_{\lambda}/2}=\mathcal{IC}(V^{\lambda}/G^{\lambda})$ be the intersection complex monodromic mixed Hodge module of the smooth stack $V^{\lambda}/G^{\lambda}$. We refer to [davison2020cohomological] for the necessary background regarding monodromic mixed Hodge modules. Alternatively, for the purposes of this paper, the reader may prefer to consider instead constructible sheaves. The Tate twist $-\otimes\mathscr{L}$ is then replaced by the shift $[-2]$.
By smoothness of $q_{\lambda}$ (Lemma 2.1), we have $q_{\lambda}^{*}\cong q_{\lambda}^{!}\otimes\mathscr{L}^{r_{\lambda}}$ and thus we obtain
$q_{\lambda}^{*}\mathbb{D}\mathbf{Q}_{V^{\lambda}/G^{\lambda}}^{\mathrm{vir}}\cong\mathbb{D}\mathbf{Q}_{V^{\lambda\geqslant 0}/P_{\lambda}}\otimes\mathscr{L}^{r_{\lambda}+d_{\lambda}/2}$
which by adjunction provides us with the map
(2.1) $\mathbb{D}\mathbf{Q}_{V^{\lambda}/G^{\lambda}}^{\mathrm{vir}}\rightarrow(q_{\lambda})_{*}\mathbb{D}\mathbf{Q}_{V^{\lambda\geqslant 0}/P_{\lambda}}\otimes\mathscr{L}^{r_{\lambda}+d_{\lambda}/2}.$
Moreover, the map $p_{\lambda}$ is proper (Lemma 2.1), and so by dualizing the adjunction map
$\mathbf{Q}_{V/G}\rightarrow(p_{\lambda})_{*}\mathbf{Q}_{V^{\lambda\geqslant 0}/P_{\lambda}},$
and using $\mathbb{D}(p_{\lambda})_{*}=(p_{\lambda})_{!}\mathbb{D}$ and $(p_{\lambda})_{*}=(p_{\lambda})_{!}$, we obtain
(2.2) $(p_{\lambda})_{*}\mathbb{D}\mathbf{Q}_{V^{\lambda\geqslant 0}/P_{\lambda}}\rightarrow\mathbb{D}\mathbf{Q}_{V/G}.$
We eventually obtain the map
(2.3) $\mathrm{Ind}_{\lambda}\colon(\imath_{\lambda})_{*}(\pi_{\lambda})_{*}\mathbb{D}\mathbf{Q}_{V^{\lambda}/G^{\lambda}}^{\mathrm{vir}}\rightarrow\pi_{*}\mathbb{D}\mathbf{Q}_{V/G}\otimes\mathscr{L}^{r_{\lambda}+d_{\lambda}/2}$
by composing $(\imath_{\lambda})_{*}(\pi_{\lambda})_{*}$ applied to (2.1) with $\pi_{*}$ applied to (2.2) twisted by $\mathscr{L}^{r_{\lambda}+d_{\lambda}/2}$.
Lemma 2.5.
Let $G$ be a symmetric representation of $G$. The induction morphism $\mathrm{Ind}_{\lambda}\colon\mathrm{H}^{*+d_{\lambda}}(V^{\lambda}/G^{\lambda})\rightarrow\mathrm{H}^{*+d}(V/G)$ preserves cohomological degrees.
Proof.
This is straightforward since the induction at the level of cohomology vector spaces is given by taking the derived global section of the morphism of complexes (2.3) (the sheafified induction) and by Lemma 2.4, $d/2=r_{\lambda}+d_{\lambda}/2$ and so in the symmetric case, the right-hand-side of (2.3) is precisely $\pi_{*}\mathbb{D}\mathbf{Q}_{V/G}^{\mathrm{vir}}$. Moreover, $\mathbb{D}\mathbf{Q}_{V/G}^{\mathrm{vir}}\cong\mathbf{Q}_{V/G}^{\mathrm{vir}}$ and $\mathbb{D}\mathbf{Q}_{V^{\lambda}/G^{\lambda}}^{\mathrm{vir}}\cong\mathbf{Q}_{V^{\lambda}/G^{\lambda}}^{\mathrm{vir}}$ since $V/G$ and $V^{\lambda}/G^{\lambda}$ are smooth and the chosen shifts make the considered constant sheaves Verdier self-dual. ∎
2.3. Augmentation
Let $\lambda\in X_{*}(T)$. We defined a representation $V^{\lambda}$ of $G^{\lambda}$ in §2.1. We can see $T$ as a maximal torus of $G^{\lambda}$. If $\mu\in X_{*}(T)$, we have $(V^{\lambda})^{\mu}=V^{\lambda}\cap V^{\mu}$, $(G^{\lambda})^{\mu}=G^{\lambda}\cap G^{\mu}$, $(V^{\lambda})^{\mu\geqslant 0}$ and $P_{\mu,\lambda}\coloneqq P_{\mu}\cap G^{\lambda}$.
In this case, the induction diagram is
[…]
The definition of the induction map associated with this datum is
$\mathrm{Ind}_{\mu,\lambda}\colon\mathrm{H}^{*}((V^{\lambda})^{\mu}/(G^{\lambda})^{\mu})\rightarrow\mathrm{H}^{*}(V^{\lambda}/G^{\lambda}).$
If $\mu\preceq\lambda$, then $(V^{\lambda})^{\mu}=V^{\mu}$ and $(G^{\lambda})^{\mu}=G^{\mu}$, so that the induction $\mathrm{Ind}_{\mu,\lambda}$ is
$\mathrm{Ind}_{\mu,\lambda}\colon\mathrm{H}^{*}(V^{\mu}/G^{\mu})\rightarrow\mathrm{H}^{*}(V^{\lambda}/G^{\lambda}).$
Lemma 2.6.
For any $\lambda,\mu\in X_{*}(T)$, one can find $\nu\in X_{*}(T)$ such that $V^{\nu}=V^{\mu}\cap V^{\lambda}$, $G^{\nu}=G^{\mu}\cap G^{\lambda}$, $(V^{\lambda})^{\nu\geqslant 0}=(V^{\lambda})^{\mu\geqslant 0}$ and $P_{\mu,\lambda}=P_{\nu,\lambda}$.
Proof.
The cocharacters $\lambda,\mu$ define a morphism
$\lambda\times\mu\colon\mathbf{G}_{\mathrm{m}}^{\times 2}\rightarrow T.$
We obtain $\nu$ by composing a general one-parameter subgroup $\mathbf{G}_{\mathrm{m}}\rightarrow\mathbf{G}_{\mathrm{m}}^{\times 2}$, $t\mapsto(t^{a},t^{b})$ with $\lambda\times\mu$. This guarantees the first two equalities of the Lemma. To guarantee the last two, it suffices to take a general one-parameter subgroup with the constraint $b>0$. ∎
2.4. Associativity
Proposition 2.7.
Let $\lambda,\mu,\nu\in\mathrm{Hom}(\mathbf{G}_{\mathrm{m}},T)$. Then,
$\mathrm{Ind}_{\lambda,\nu}=\mathrm{Ind}_{\mu,\nu}\circ\mathrm{Ind}_{\lambda,\mu}.$
Proof.
By Lemma 2.6, we may assume that $\lambda\preceq\mu\preceq\nu$. This assumption only simplifies the notations in diagram (2.4). The equality then classically follows by base-change in the commutative diagram
[…]
2.5. Explicit formula for the parabolic induction
For $\lambda\in X_{*}(T)$ we define the induction kernel
$k_{\lambda}\coloneqq\frac{\prod_{\alpha\in\mathcal{W}^{\lambda>0}(V)}\alpha}{\prod_{\alpha\in\mathcal{W}^{\lambda>0}(\mathfrak{g})}\alpha}\in\mathrm{Frac}(\mathrm{H}^{*}(\mathrm{pt}/T)).$
We let $W^{\lambda}\coloneqq N_{G^{\lambda}}(T)/T$ be the Weyl group of $G^{\lambda}$.
Lemma 2.8.
For any $\lambda\in X_{*}(T)$, $k_{\lambda}$ is $W^{\lambda}$-invariant.
Proof.
It suffices to prove that $W^{\lambda}$ send $\mathcal{W}^{\lambda>0}(V)\subseteq\mathcal{W}(W)$ to itself, and similarly for $\mathcal{W}^{\lambda>0}(\mathfrak{g})$. This comes from the invariance of the pairing between characters and cocharacters: for any $w\in W$, $\langle w\cdot\lambda,w\cdot\alpha\rangle=\langle\lambda,\alpha\rangle$ and if $w\in W^{\lambda}$, $w\cdot\lambda=\lambda$. ∎
Proposition 2.9.
Let $\lambda\colon\mathbf{G}_{\mathrm{m}}\rightarrow T$ be a cocharacter and $f\in\mathrm{H}^{*}(V^{\lambda}/G^{\lambda})$. We have
$\mathrm{Ind}_{\lambda}(f)=\sum_{w\in W/W^{\lambda}}w\cdot(fk_{\lambda})=\frac{1}{\#W^{\lambda}}\sum_{w\in W}w\cdot(fk_{\lambda}).$
Proof.
The second equality follows from the $W^{\lambda}$-invariance of both $f$ (since $f\in\mathrm{H}^{*}(V^{\lambda}/T)^{W^{\lambda}}$) and $k_{\lambda}$ (Lemma 2.8). The first equality is a computation of Euler class exactly as in the proof of [kontsevich2011cohomological, Theorem 2] in the case of quiver representations. ∎
More generally, let $\lambda,\mu\in X_{*}(T)$. We let
$k_{\lambda,\mu}\coloneqq\frac{\prod_{\alpha\in\mathcal{W}^{\lambda>0}(V^{\mu})}\alpha}{\prod_{\alpha\in\mathcal{W}^{\lambda>0}(\mathfrak{g}^{\mu})}\alpha}.$
Proposition 2.10 (Augmentation).
For $f\in\mathrm{H}^{*}((V^{\mu})^{\lambda}/(G^{\mu})^{\lambda})$, we have
$\mathrm{Ind}_{\lambda,\mu}(f)=\sum_{w\in W^{\mu}/(W^{\lambda}\cap W^{\mu})}w(fk_{\lambda,\mu}).$
Proof.
This follows immediately from Proposition 2.9 applied to $V^{\mu}$ as a representation of $G^{\mu}$ and the cocharacter $\lambda$ of $T$, seen as a maximal torus of $G^{\mu}$, and the identification of the Weyl group of $(G^{\mu})^{\lambda}=G^{\lambda}\cap G^{\mu}$ with $W^{\lambda}\cap W^{\mu}$. ∎
2.6. Parabolic induction for symmetric representations
Proposition 2.11.
Let $V$ be a symmetric representation of a reductive group $G$ and $\lambda\in X_{*}(T)$ a cocharacter of the maximal torus of $T$. Then, the representation $V^{\lambda}$ of $G^{\lambda}$ is also symmetric.
Proof.
It is immediate to check that $\mathcal{W}^{\prime}(V^{\lambda})=\{\alpha\in\mathcal{W}^{\prime}(V)\mid\langle\lambda,\alpha\rangle=0\}$ and for any $\alpha\in\mathcal{W}^{\prime}(V^{\lambda})$, we can identify the weight spaces $V^{\lambda}_{\alpha}=V_{\alpha}$. Therefore, the symmetry of $V$ implies that of $V^{\lambda}$. ∎
Recall the order relation on $X_{*}(T)$ defined in the introduction
[…]
for any $\lambda,\mu\in X_{*}(T)$.
We obtain an equivalence relation on the set $X_{*}(T)$
[…]
We let $\mathscr{P}_{V}\coloneqq X_{*}(T)/\sim$ be the quotient of $X_{*}(T)$ by this equivalence relation. If $\lambda\in X_{*}(T)$, $\overline{\lambda}$ denotes its class in $\mathscr{P}_{V}$.
It has an induced order relation still denoted by $\preceq$:
$\overline{\lambda}\preceq\overline{\mu}\iff\lambda\preceq\mu.$
It is immediate that this is well-defined, i.e. the definition does not depend on the representatives $\lambda,\mu$ of $\overline{\lambda},\overline{\mu}$.
Lemma 2.12.
Let $V$ be a symmetric representation of a reductive group $G$. If $\lambda,\mu\in X_{*}(T)$ are such that $\lambda\sim\mu$, then #$\#\mathcal{W}^{\lambda>0}(V)=\#\mathcal{W}^{\mu>0}(V)=\#\mathcal{W}^{\lambda<0}(V)=\#\mathcal{W}^{\mu<0}(V)$ and #$\#\mathcal{W}^{\lambda>0}(\mathfrak{g})=\#\mathcal{W}^{\mu>0}(\mathfrak{g})=\#\mathcal{W}^{\lambda<0}(\mathfrak{g})=\#\mathcal{W}^{\mu<0}(\mathfrak{g})$.
Proof.
We prove the statement for $V$. The statement for $\mathfrak{g}$ is obtained by specialising $V=\mathfrak{g}$ since $\mathfrak{g}$ under the adjoint action of $G$ is a symmetric representation of $G$.
Since $V$ is symmetric, we have #$\#\mathcal{W}^{\lambda>0}(V)=\#\mathcal{W}^{\lambda<0}(V)$ and moreover, #$\#\mathcal{W}(V)=\#\mathcal{W}^{\lambda>0}(V)+\#\mathcal{W}^{\lambda<0}(V)+\mathcal{W}(V^{\lambda})$. Since we have the same equalities with $\lambda$ replaced by $\mu$ and $\mathcal{W}^{\lambda=0}(V)=\mathcal{W}^{\mu=0}(V)$, we obtain #$\#\mathcal{W}^{\lambda>0}(V)=\#\mathcal{W}^{\mu>0}(V)$. The other equalities follow. ∎
Lemma 2.13.
If $\lambda,\mu\in X_{*}(T)$ are such that $\lambda\sim\mu$, then $k_{\lambda}=\pm k_{\mu}$.
Proof.
If $\lambda\sim\mu$, we have $\mathcal{W}^{\lambda>0}(V)\sqcup\mathcal{W}^{\lambda<0}(V)=\mathcal{W}^{>\mu}(V)\sqcup\mathcal{W}^{<\mu}(V)$ and $\mathcal{W}^{\lambda>0}(\mathfrak{g})\sqcup\mathcal{W}^{\lambda<0}(\mathfrak{g})=\mathcal{W}^{>\mu}(\mathfrak{g})\sqcup\mathcal{W}^{<\mu}(\mathfrak{g})$ and since $V$ and $\mathfrak{g}$ are symmetric representations of $G$, $\mathcal{W}^{\lambda>0}(V)=-\mathcal{W}^{\lambda<0}(V)$, $\mathcal{W}^{>\mu}(V)=-\mathcal{W}^{<\mu}(V)$ and similarly for $\mathfrak{g}$. Therefore,
$(-1)^{\#\mathcal{W}^{\lambda>0}(V)-\#\mathcal{W}^{\lambda>0}(\mathfrak{g})}k_{\lambda}^{2}=\frac{\prod_{\alpha\in\mathcal{W}^{\lambda>0}(V)\sqcup\mathcal{W}^{\lambda<0}(V)}\alpha}{\prod_{\alpha\in\mathcal{W}^{\lambda>0}(\mathfrak{g})\sqcup\mathcal{W}^{\lambda<0}(\mathfrak{g})}\alpha}=\frac{\prod_{\alpha\in\mathcal{W}^{\mu>0}(V)\sqcup\mathcal{W}^{\mu<0}(V)}\alpha}{\prod_{\alpha\in\mathcal{W}^{\mu>0}(\mathfrak{g})\sqcup\mathcal{W}^{\mu<0}(\mathfrak{g})}\alpha}=(-1)^{\#\mathcal{W}^{\mu>0}(V)-\#\mathcal{W}^{\mu>0}(\mathfrak{g})}k_{\mu}^{2}.$
Since by Lemma 2.12 we have #$\#\mathcal{W}^{\lambda>0}(V)=\#\mathcal{W}^{\lambda<0}(V)=\#\mathcal{W}^{\mu>0}(V)=\#\mathcal{W}^{\mu<0}(V)$ and #$\#\mathcal{W}^{\lambda>0}(\mathfrak{g})=\#\mathcal{W}^{\lambda<0}(\mathfrak{g})=\#\mathcal{W}^{\mu>0}(\mathfrak{g})=\#\mathcal{W}^{\mu<0}(\mathfrak{g})$, we have $k_{\lambda}^{2}=k_{\mu}^{2}$ and therefore, $k_{\lambda}=\pm k_{\mu}$. ∎
Lemma 2.14.
We assume that $V$ is a symmetric representation of $G$. Then, for any $\lambda,\mu\in X_{*}(T)$ such that $\overline{\lambda}=\overline{\mu}$, the induction maps $\mathrm{Ind}_{\lambda}$ and $\mathrm{Ind}_{\mu}$ differ by a sign. In particular, the images of $\mathrm{Ind}_{\lambda}$ and $\mathrm{Ind}_{\mu}$ inside $\mathrm{H}^{*+d}(V/G)$ coincide.
Proof.
We use the explicit formula for the induction (Proposition 2.9). Let $f\in\mathrm{H}^{*}(V^{\lambda}/G^{\lambda})=\mathrm{H}^{*}(V^{\mu}/G^{\mu})$. We have
$\mathrm{Ind}_{\lambda}(f)=\sum_{w\in W/W^{\lambda}}w\cdot(fk_{\lambda}),\quad\mathrm{Ind}_{\mu}(f)=\sum_{w\in W/W^{\mu}}w\cdot(fk_{\mu}).$
Moreover, $G^{\lambda}=G^{\mu}$, $W^{\lambda}=W^{\mu}$ and by Lemma 2.13, $k_{\mu}=\pm k_{\lambda}$. This concludes. ∎
The $W$-action on $X_{*}(T)$ induces a $W$-action on $\mathscr{P}_{V}$. For $\lambda\in X_{*}(T)$, we let $\widetilde{\lambda}\in\mathscr{P}_{V}/W$ be its $W$-orbit.
Lemma 2.15.
Let $\lambda,\mu\in X_{*}(T)$ be such that $\widetilde{\lambda}=\widetilde{\mu}$. Then, the images of $\mathrm{Ind}_{\lambda}$ and $\mathrm{Ind}_{\mu}$ coincide.
Proof.
By Lemma 2.14, we may assume that for some $w\in W$, $\mu=w\cdot\lambda$. Then, $\mathcal{W}^{\mu>0}(V)=w\cdot\mathcal{W}^{\lambda>0}(V)$ and $\mathcal{W}^{\mu>0}(\mathfrak{g})=w\cdot\mathcal{W}^{\lambda>0}(\mathfrak{g})$. Therefore, $w(k_{\lambda})=k_{\mu}$. Using the explicit formula for the induction product (Proposition 2.9), we obtain
$\mathrm{Ind}_{\mu}(f)=\mathrm{Ind}_{\lambda}(w^{-1}(f)),$
and therefore, since $w^{-1}$ induces an isomorphism $\mathrm{H}^{*}(V^{\mu}/G^{\mu})=\mathrm{H}^{*}(V^{w\cdot\lambda}/G^{w\cdot\lambda})\cong\mathrm{H}^{*}(V^{\lambda}/G^{\lambda})$, the images of $\mathrm{Ind}_{\lambda}$ and $\mathrm{Ind}_{\mu}$ coincide. ∎
3. Tautological classes
Let $G$ be a reductive group and $V$ a representation of $G$. We let $\pi\colon V/G\rightarrow V/\!\!/G$ be the good moduli space map. We let $\kappa\colon V/\!\!/G\rightarrow\mathrm{pt}$ be the projection to the point.
Then, for any line bundle $\mathcal{L}$ over $V/G$, we have an operation $\pi_{*}\mathbf{Q}_{V/L}\rightarrow\pi_{*}\mathbf{Q}_{V/L}$ of multiplication by the first Chern class $c_{1}(\mathcal{L})$. More precisely, we have the following lemma.
Lemma 3.1.
We have a morphism of graded rings
[…]
such that the composition $\kappa_{*}\circ a$ where $\kappa_{*}\colon\mathrm{RHom}(\pi_{*}\mathbf{Q}_{V/G},\pi_{*}\mathbf{Q}_{V/G})\rightarrow\mathrm{RHom}(\kappa_{*}\pi_{*}\mathbf{Q}_{V/G},\kappa_{*}\pi_{*}\mathbf{Q}_{V/G})\cong\mathrm{RHom}(\mathrm{H}^{*}(\mathrm{pt}/G),\mathrm{H}^{*}(\mathrm{pt}/G))$, send $a\in\mathrm{H}^{*}(\mathrm{pt}/G)$ to the endomorphism of $\mathrm{H}^{*}(\mathrm{pt}/G)$ given by the cup product with $a$.
Proof.
We have a canonical adjunction morphism
$\mathrm{RHom}(\pi_{*}\mathbf{Q}_{V/G},\pi_{*}\mathbf{Q}_{V/G})\cong\mathrm{RHom}(\pi^{*}\pi_{*}\mathbf{Q}_{V/G},\mathbf{Q}_{V/G}).$
By precomposing with the counit map $\pi^{*}\pi_{*}\mathbf{Q}_{V/G}\rightarrow\mathbf{Q}_{V/G}$, we obtain the morphism
$\mathrm{RHom}(\mathbf{Q}_{V/G},\mathbf{Q}_{V/G})\cong\mathrm{H}^{*}(V/G)\cong\mathrm{H}^{*}(\mathrm{pt}/G)\rightarrow\mathrm{RHom}(\pi_{*}\mathbf{Q}_{V/G},\pi_{*}\mathbf{Q}_{V/G}).$
The last statement of the theorem follows from the functoriality of adjunctions. ∎
We now state and prove an elementary lemma regarding Cartan subalgebras of reductive Lie algebras.
Lemma 3.2.
Let $\mathfrak{g}$ be a reductive Lie algebra, $\mathfrak{h}\subseteq\mathfrak{g}$ an ideal and $\mathfrak{t}$ a Cartan subalgebra. Then, $\mathfrak{h}\cap\mathfrak{t}$ is a Cartan subalgebra of $\mathfrak{h}$.
Proof.
Let $\pi\colon\mathfrak{g}\rightarrow\mathfrak{g}/\mathfrak{h}$ be the projection. Since $\mathfrak{g}$ is reductive, we have a (noncanonical) decomposition $\mathfrak{g}\cong\mathfrak{h}\oplus(\mathfrak{g}/\mathfrak{h})$. If $\mathfrak{t}^{\prime}$ is a maximal torus of $\mathfrak{h}$ and $\mathfrak{t}^{\prime\prime}$ is a maximal torus of $\mathfrak{g}/\mathfrak{h}$, then $\mathfrak{t}^{\prime}\oplus\mathfrak{t}^{\prime\prime}$ is a maximal torus of $\mathfrak{g}$. Moreover, $\pi(\mathfrak{t})\cong\mathfrak{t}/\mathfrak{t}\cap\mathfrak{h}$ is an Abelian subalgebra of $\mathfrak{g}/\mathfrak{h}$ and $\mathfrak{t}\cap\mathfrak{h}$ is an Abelian subalgebra of $\mathfrak{h}$. By maximality of $\dim\mathfrak{t}^{\prime}$, we have
$\dim\pi(\mathfrak{t})=\dim\mathfrak{t}-\dim(\mathfrak{t}\cap\mathfrak{h})\geqslant\dim\mathfrak{t}-\dim\mathfrak{t}^{\prime}=\dim\mathfrak{t}^{\prime\prime}$
By maximality of $\dim\mathfrak{t}^{\prime\prime}$, we have equality in this chain of inequalities. This means that $\dim(\mathfrak{t}\cap\mathfrak{h})=\dim\mathfrak{t}^{\prime}$ and therefore, $\mathfrak{t}\cap\mathfrak{h}$ is a Cartan subalgebra of $\mathfrak{h}$. ∎
We have the group version of Lemma 3.2.
Lemma 3.3.
Let $G$ be a reductive group, $H\subseteq G$ a normal subgroup and $T\subseteq G$ a maximal torus. Then, the neutral component of $T\cap H$ is a maximal torus of $H$.
Proof.
This follows from Lemma 3.2 by taking the Lie algebras of all groups appearing. ∎
Example 3.4.
It is necessary to consider the neutral component of $T\cap H$ as shown by the elementary case of $H=\{\pm 1\}\subseteq\mathrm{SL}_{2}(\mathbf{C})$, and where $T\subseteq\mathrm{SL}_{2}(\mathbf{C})$ is the standard torus.
Proposition 3.5.
Let $G$ be a reductive group. For any normal subgroup $H\subseteq G$ acting trivially on $V$, we have (non-canonical) actions of $\mathrm{H}^{*}(\mathrm{pt}/H)$ on $\mathrm{H}^{*}(\mathrm{pt}/G)\cong\mathrm{H}^{*}(V/G)$ and $\pi_{*}\mathbf{Q}_{V/G}$.
Proof.
By choosing a decomposition $\mathrm{Lie}(G)\cong\mathrm{Lie}(H)\oplus\mathrm{Lie}(G/H)$, we obtain a decomposition $\mathrm{H}^{*}(\mathrm{pt}/G)\cong\mathrm{H}^{*}(\mathrm{pt}/(G/H))\otimes\mathrm{H}^{*}(\mathrm{pt}/H)$. At the sheaf level, the action of the lemma is obtained by pre-composing the morphism $a$ of Lemma 3.1 with the inclusion $\mathrm{H}^{*}(\mathrm{pt}/H)\rightarrow\mathrm{H}^{*}(\mathrm{pt}/G)$. ∎
Lemma 3.6.
Let $G$ be an algebraic group and $H\subseteq G$ a normal subgroup. Then, we have a natural inclusion $\mathrm{H}^{*}(\mathrm{pt}/(G/H))\rightarrow\mathrm{H}^{*}(\mathrm{pt}/G)$.
Proof.
The map $\mathrm{H}^{*}(\mathrm{pt}/(G/H))\rightarrow\mathrm{H}^{*}(\mathrm{pt}/G)$ is the pullback in cohomology for the morphism of stacks $\mathrm{pt}/(G/H)\rightarrow\mathrm{pt}/G$ induced by the surjective morphism $G\rightarrow G/H$. ∎
Proposition 3.7.
Let $G$ be an algebraic group, and $V$ a representation of $G$. Let $H$ be a normal subgroup of $G$ acting trivially on $V$. Then, the morphism
$\mathrm{H}^{*}(\mathrm{pt}/H)\otimes\mathrm{H}^{*}(V/(G/H))\rightarrow\mathrm{H}^{*}(V/G)$
obtained by combining the natural pullback map $\mathrm{H}^{*}(V/(G/H))\rightarrow\mathrm{H}^{*}(V/G)$ coming from the morphism of stacks $V/G\rightarrow V/(G/H)$ (Lemma 3.6) with the $\mathrm{H}^{*}(\mathrm{pt}/H)$-action on $\mathrm{H}^{*}(V/G)$ (Proposition 3.5) is an isomorphism.
Proof.
We may assume that $V=\mathrm{pt}$, that $G$ is connected and by quotienting out by the unipotent radical of $G$, that $G$ is reductive. We let $\mathfrak{g}=\mathrm{Lie}(G)$ and $\mathfrak{h}=\mathrm{Lie}(H)$. We let $\mathfrak{t}\subseteq\mathfrak{g}$ be a Cartan subalgebra. We have $\mathfrak{g}\cong\mathfrak{h}\oplus(\mathfrak{g}/\mathfrak{h})$ and therefore a decomposition of the Weyl group $W_{\mathfrak{g}}\cong W_{\mathfrak{h}}\times W_{\mathfrak{g}/\mathfrak{h}}$. We let $\pi\colon\mathfrak{g}\rightarrow\mathfrak{g}/\mathfrak{h}$ be the projection. The morphism of the proposition can then be identified with the isomorphism
$\mathrm{Sym}((\mathfrak{t}\cap\mathfrak{h})^{*})^{W_{\mathfrak{h}}}\otimes\mathrm{Sym}(\pi(\mathfrak{t})^{*})^{W_{\mathfrak{g}/\mathfrak{h}}}\rightarrow\mathrm{Sym}(\mathfrak{t}^{*})^{W}.$
4. Cohomological integrality for symmetric representations of reductive groups
We fix a symmetric representation $V$ of a connected reductive group $G$.
4.1. Weyl groups
Let $\lambda\colon\mathbf{G}_{\mathrm{m}}\rightarrow T$. We let $G^{\lambda}\subseteq G$ be the corresponding Levi subgroup and $G_{\lambda}$ the intersection of the kernel of the action of $G^{\lambda}$ on $V^{\lambda}$ and the center of $G^{\lambda}$. We let $W^{\lambda}=N_{G^{\lambda}}(T)$ be the Weyl group of $G^{\lambda}$ and $W_{\lambda}\coloneqq\{w\in W\mid V^{w\cdot\lambda}=V^{\lambda}\text{ and }G^{w\cdot\lambda}=G^{\lambda}\}=\{w\in W\mid w\cdot\overline{\lambda}=\overline{\lambda}\}$.
Lemma 4.1.
$W_{\lambda}$ is a subgroup of $W$.
Proof.
We let $\dot{w}\in N_{G}(T)$ be a lift of $w\in W$. Then, $V^{w\cdot\lambda}=\dot{w}V^{\lambda}$. The statement of the lemma follows then immediately from the fact that if $w,w^{\prime}\in W$, and $\dot{w},\dot{w^{\prime}}\in N_{G}(T)$ are respective lifts, then, $\dot{w}\dot{w^{\prime}}$ is a lift of $ww^{\prime}$ and $\dot{w}^{-1}$ a lift of $w^{-1}$. ∎
Lemma 4.2.
We have natural inclusions $W^{\lambda}\subseteq W_{\lambda}\subseteq W$.
Proof.
The inclusion $W_{\lambda}\subseteq W$ holds by definition of $W_{\lambda}$ (see Lemma 4.1). The inclusion $W^{\lambda}\subseteq W$ comes from the inclusion $N_{G^{\lambda}}(T)\subseteq N_{G}(T)$ since $G$ and $G^{\lambda}$ share the same maximal torus $T$.
Let $w\in W^{\lambda}$ and $\dot{w}\in N_{G^{\lambda}}(T)$ a lift of $w$. Then, $\dot{w}G^{\lambda}\dot{w}^{-1}=G^{w\cdot\lambda}=G^{\lambda}$ and $\dot{w}V^{\lambda}=V^{w\cdot\lambda}=V^{\lambda}$ and so $w\in W_{\lambda}$. ∎
Lemma 4.3.
The group $W_{\lambda}$ normalises the group $W^{\lambda}$ inside $W$.
Proof.
It suffices to check that if $w\in W_{\lambda}$ and $w^{\prime}\in W^{\lambda}$, and $\dot{w}\in N_{G}(T)$, $\dot{w^{\prime}}\in N_{G^{\lambda}}(T)$ are any lifts, then $\dot{w}\dot{w^{\prime}}\dot{w}^{-1}\in G^{\lambda}$. This comes from the fact that $\dot{w}$ and $\lambda$ commute and $\lambda$ fixes $G^{\lambda}$ and so $\dot{w}\dot{w^{\prime}}\dot{w}^{-1}$ is fixed by $\lambda$. ∎
Corollary 4.4.
The group $W_{\lambda}$ acts on $\mathrm{H}^{*}(\mathrm{pt}/T)^{W^{\lambda}}\cong\mathrm{H}^{*}(\mathrm{pt}/G^{\lambda})$.
Proof.
As the Weyl group $W$ of $G$ acts on $\mathrm{H}^{*}(\mathrm{pt}/T)$. We just have to check that the induced action of the subgroup $W_{\lambda}$ preserves $W^{\lambda}$-invariants. This is a consequence of Lemma 4.3. ∎
Lemma 4.5.
The group $W_{\lambda}$ acts naturally on $\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$.
Proof.
By definition, $G_{\lambda}$ is preserved by $W_{\lambda}$. ∎
We let $\overline{G^{\lambda}}\coloneqq G^{\lambda}/G_{\lambda}$. Since $G_{\lambda}$ is a normal subgroup of $G^{\lambda}$, $\overline{G^{\lambda}}$ is a reductive group. Since $G_{\lambda}$ is contained in the center of $G^{\lambda}$, the Weyl group of $\overline{G^{\lambda}}$ is isomorphic to $W$.
Lemma 4.6.
The group $W_{\lambda}$ acts naturally on $\mathrm{H}^{*}(V^{\lambda}/\overline{G^{\lambda}})$.
Proof.
We have $\mathrm{H}^{*}(V^{\lambda}/\overline{G^{\lambda}})\cong\mathrm{H}^{*}(\mathrm{pt}/(T/G_{\lambda}))^{W^{\lambda}}$.
The restriction to $W_{\lambda}$ of the $W$ action on $T$ preserves $G_{\lambda}$ since $G_{\lambda}$ is in the center of $G^{\lambda}$. Therefore, we obtain a $W_{\lambda}$-action on $T/G_{\lambda}$. It induces a $W_{\lambda}$-action on $\mathrm{H}^{*}(\mathrm{pt}/(T/G_{\lambda}))$. The fact that $W_{\lambda}$ normalises $W^{\lambda}$ (Lemma 4.3) implies that $\mathrm{H}^{*}(\mathrm{pt}/(T/G_{\lambda}))^{W^{\lambda}}$ is preserved by the $W_{\lambda}$-action. ∎
By combining Corollary 4.4, Lemma 4.5 and Lemma 4.6, we obtain the following lemma.
Lemma 4.7.
The $W_{\lambda}$-action on $\mathrm{H}^{*}(\mathrm{pt}/G^{\lambda})\cong\mathrm{H}^{*}(V^{\lambda}/G^{\lambda})$ obtained in Corollary 4.4 and the $W_{\lambda}$-action on $\mathrm{H}^{*}(V^{\lambda}/\overline{G^{\lambda}})\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$ obtained by the tensor product of the actions given in Lemmas 4.5 and 4.6 coincide via the isomorphism given in Proposition 3.7.∎
Proposition 4.8.
There exists a character $\varepsilon_{V,\lambda}\colon W_{\lambda}\rightarrow\{\pm 1\}$ such that for any $w\in W_{\lambda}$, $w(k_{\lambda})=\varepsilon_{V,\lambda}(w)k_{\lambda}$.
Proof.
It suffices to prove the existence and unicity of the sign $\varepsilon_{V,\lambda}(w)\in\{\pm 1\}$ satisfying the equality of the proposition, for any $w\in W_{\lambda}$. The multiplicativity then follows from the unicity of this sign.
The set of weights $\mathcal{W}^{\lambda=0}(V)=\mathcal{W}(V^{\lambda})$ only depends on $\overline{\lambda}$ and so is stable under the action of $W_{\lambda}$. Therefore, $\mathcal{W}^{\lambda>0}(V)\sqcup\mathcal{W}^{\lambda<0}(V)=\mathcal{W}(V)\smallsetminus\mathcal{W}^{\lambda=0}(V)$ is stable under $W_{\lambda}$. Also, by symmetry, $\mathcal{W}^{\lambda>0}(V)=-\mathcal{W}^{\lambda<0}(V)$. Therefore,
$\prod_{\alpha\in\mathcal{W}^{\lambda>0}(V)\sqcup\mathcal{W}^{\lambda<0}(V)}\alpha=(-1)^{\#\mathcal{W}^{\lambda>0}(V)}\prod_{\alpha\in\mathcal{W}^{\lambda>0}(V)}\alpha^{2}.$
By applying this equality to $\mathfrak{g}$ under the adjoint action, seen as a symmetric representation of $G$, we obtain
$\prod_{\alpha\in\mathcal{W}^{\lambda>0}(\mathfrak{g})\sqcup\mathcal{W}^{\lambda<0}(\mathfrak{g})}\alpha=(-1)^{\#\mathcal{W}^{\lambda>0}(\mathfrak{g})}\prod_{\alpha\in\mathcal{W}^{\lambda>0}(\mathfrak{g})}\alpha^{2}.$
Therefore,
$(-1)^{\#\mathcal{W}^{\lambda>0}(V)-\#\mathcal{W}^{\lambda>0}(\mathfrak{g})}k_{\lambda}^{2}=\frac{\prod_{\alpha\in\mathcal{W}^{\lambda>0}(V)\sqcup\mathcal{W}^{\lambda<0}(V)}\alpha}{\prod_{\alpha\in\mathcal{W}^{\lambda>0}(\mathfrak{g})\sqcup\mathcal{W}^{\lambda<0}(\mathfrak{g})}\alpha}.$
By applying $w\in W_{\lambda}$, we obtain $w(k_{\lambda}^{2})=k_{\lambda}^{2}$. Therefore, $w(k_{\lambda})=\pm k_{\lambda}$, which proves the existence of the sign. The unicity is straighforward. ∎
Proposition 4.9.
Let $\lambda\in X_{*}(T)$, $w\in W_{\lambda}$ and $f\in\mathcal{H}_{\lambda}$. Then, $\mathrm{Ind}_{\lambda}(w\cdot f)=\varepsilon(w)\mathrm{Ind}_{\lambda}(f)$.
Proof.
By Proposition 2.9, we have
$\mathrm{Ind}_{\lambda}(w\cdot f)$ $=\frac{1}{\#W^{\lambda}}\sum_{w^{\prime}\in W}w^{\prime}(wf)w^{\prime}(k_{\lambda})$
$=\frac{1}{\#W^{\lambda}}\sum_{w^{\prime\prime}\in W}(w^{\prime\prime}f)w^{\prime\prime}(w^{-1}k_{\lambda})$
[…]
which concludes. ∎
We now give an elementary result from the representation theory of finite groups.
Proposition 4.10.
Let $G$ be a finite group $\chi\colon G\rightarrow\mathbf{C}^{*}$ a character of $G$ and $V$ a finite-dimensional representation of $G$. Then, the formula
$p_{\chi}(v)=\frac{1}{\#G}\sum_{g\in G}\chi(g)^{-1}g\cdot v$
is the projector onto the $\chi$-isotypic component $V^{\chi}$ of $V$.
Proof.
There are essentially three things to check:
  1. (1)
    $p_{\chi}$ is an involution: $p_{\chi}\circ p_{\chi}=p_{\chi}$,
  2. (2)
    $p_{\chi}(f)=f$ for $f\in V^{\chi}$,
  3. (3)
    $\mathrm{im}(p_{\chi})\subseteq V^{\chi}$,
which are routine. ∎
4.2. Proof of the cohomological integrality theorem
In the rest of this section, we prove Theorem 1.1.
For $\lambda\in X_{*}(T)$, we define the vector space
$\mathcal{H}_{\lambda}\coloneqq\mathrm{H}^{*+d_{\lambda}}(V^{\lambda}/G^{\lambda})$
By §2.3 We have induction maps $\mathrm{Ind}_{\lambda,\mu}\colon\mathcal{H}_{\lambda}\rightarrow\mathcal{H}_{\mu}$ for $\lambda,\mu\in X_{*}(T)$.
For $\lambda\in X_{*}(T)$, we choose a splitting $\mathfrak{g}^{\lambda}=\mathfrak{g}_{\lambda}\oplus(\mathfrak{g}^{\lambda}/\mathfrak{g}_{\lambda})$. It induces a splitting of the maximal torus $\mathfrak{t}=\mathfrak{g}_{\lambda}\oplus(\mathfrak{t}/\mathfrak{g}_{\lambda})$ (Lemma 3.2). We let $\mathcal{H}_{\lambda}^{\mathrm{prim}}\coloneqq\mathrm{H}^{*}(V^{\lambda}/(G^{\lambda}/G_{\lambda}))$, so that $\mathcal{H}_{\lambda}\cong\mathcal{H}_{\lambda}^{\mathrm{prim}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$ (Proposition 3.7). We let $\mathcal{A}_{\lambda}\coloneqq\mathrm{H}^{*}(V^{\lambda}/T)$ and $\mathcal{A}_{\lambda}^{\mathrm{prim}}\coloneqq\mathrm{H}^{*}(V^{\lambda}/(T/G_{\lambda}))$ so that $\mathcal{A}_{\lambda}\cong\mathcal{A}_{\lambda}^{\mathrm{prim}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$, $\mathcal{H}_{\lambda}=\mathcal{A}_{\lambda}^{W^{\lambda}}$ and $\mathcal{H}_{\lambda}^{\mathrm{prim}}\cong(\mathcal{A}_{\lambda}^{\mathrm{prim}})^{W^{\lambda}}$.
For each $\lambda\in X_{*}(T)$, we let $J_{\lambda}$ be the smallest $W^{\lambda}$-stable $\mathcal{A}_{\lambda}^{\mathrm{prim}}$-submodule of $\mathcal{A}_{\lambda}^{\mathrm{prim},\mathrm{loc}}\coloneqq\mathcal{A}_{\lambda}^{\mathrm{prim}}[\prod_{\alpha\in\mathcal{W}(\mathfrak{g}^{\lambda})\smallsetminus\{0\}}\alpha^{-1}]$ containing
$k_{\mu,\lambda}=\frac{\prod_{\alpha\in\mathcal{W}^{\mu>0}(V^{\lambda})}\alpha}{\prod_{\alpha\in\mathcal{W}^{\mu>0}(\mathfrak{g}^{\lambda})}\alpha}$
for all $\mu\in X_{*}(T)$ such that $\lambda\not\preceq\mu$.
We let $\mathcal{A}_{\lambda}^{\mathrm{loc}}\coloneqq\mathcal{A}_{\lambda}[\prod_{\alpha\in\mathcal{W}(\mathfrak{g}^{\lambda})\smallsetminus\{0\}}\alpha^{-1}]\cong\mathcal{A}_{\lambda}^{\mathrm{prim},\mathrm{loc}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$ and $\mathcal{H}_{\lambda}^{\mathrm{loc}}\coloneqq(\mathcal{A}_{\lambda}^{\mathrm{loc}})^{W^{\lambda}}\cong\mathcal{H}_{\lambda}^{\mathrm{prim}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$.
Lemma 4.11.
We can characterise $J_{\lambda}$ as the smallest $W^{\lambda}$-stable $\mathcal{A}_{\lambda}^{\mathrm{prim}}$-submodule of $\mathcal{A}_{\lambda}^{\mathrm{prim}}[\prod_{\alpha\in\mathcal{W}(\mathfrak{g}^{\lambda})\smallsetminus\{0\}}\alpha^{-1}]$ containing
$k_{\mu,\lambda}=\frac{\prod_{\alpha\in\mathcal{W}^{\mu>0}(V^{\lambda})}\alpha}{\prod_{\alpha\in\mathcal{W}^{\mu>0}(\mathfrak{g}^{\lambda})}\alpha}$
for all $\mu\in X_{*}(T)$ such that $\mu\prec\lambda$.
Proof.
If $\mu\prec\lambda$, then we have $\dim V^{\mu}\leqslant\dim V^{\lambda}$ and $\dim\mathfrak{g}^{\mu}\leqslant\dim\mathfrak{g}^{\lambda}$ with at least one strict inequality. If we have $\lambda\preceq\mu$ then, we have
$\dim V^{\lambda}\leqslant\dim V^{\mu},\quad\dim\mathfrak{g}^{\lambda}\leqslant\dim\mathfrak{g}^{\mu}$
providing a contradiction. Therefore, if $\mu\prec\lambda$, then $\lambda\not\preceq\mu$ and by definition, $k_{\mu,\lambda}\in J_{\lambda}$.
Conversely, if $\lambda\not\preceq\mu$, we let $\nu\in X_{*}(T)$ be such that $V^{\nu}=V^{\lambda}\cap V^{\mu}$, $\mathfrak{g}^{\nu}=\mathfrak{g}^{\lambda}\cap\mathfrak{g}^{\mu}$, $(V^{\lambda})^{\nu\geqslant 0}=(V^{\lambda})^{\mu\geqslant 0}$ and $(\mathfrak{g}^{\lambda})^{\nu\geqslant 0}=(\mathfrak{g}^{\lambda})^{\mu\geqslant 0}$ (Lemma 2.6). We have $V^{\nu}\subseteq V^{\lambda}$ and $\mathfrak{g}^{\nu}\subseteq\mathfrak{g}^{\lambda}$ with at least one strict inclusion. Therefore, $\nu\prec\lambda$ and moreover, $k_{\nu,\lambda}=k_{\mu,\lambda}$. This proves the lemma. ∎
Lemma 4.12.
We have $J_{\lambda}^{W^{\lambda}}\subseteq\mathcal{H}_{\lambda}^{\mathrm{prim}}$.
Proof.
This comes from the fact that the induction product §2.2 is well-defined on the cohomology, without requiring localisation, even though the explicit formula (Proposition 2.9) involves rational fractions. More precisely, $J_{\lambda}^{W^{\lambda}}$ is linearly generated by the elements
$\sum_{w\in W^{\lambda}}w\cdot(fk_{\mu,\lambda})$
for $f\in\mathcal{A}_{\lambda}^{\mathrm{prim}}$ and $\mu\prec\lambda$, by Lemma 4.11. These are averages over $W^{\lambda}$ of elements of $J_{\lambda}$. We can rewrite this sum as
$\sum_{\overline{w^{\prime}}\in W^{\lambda}/W^{\mu}}w^{\prime}\cdot\left(\sum_{w\in W^{\mu}}(w\cdot f)\right)k_{\mu,\lambda}$
since $k_{\mu,\lambda}$ is $W^{\mu}$-invariant (Lemma 2.8). This is the formula for
$I\coloneqq\mathrm{Ind}_{\mu,\lambda}\left(\sum_{w\in W^{\mu}}(w\cdot f)\right).$
Since $\sum_{w\in W^{\mu}}(w\cdot f)$ is polynomial, its induction $I$ is also polynomial. Therefore, it belongs to $\mathcal{H}_{\lambda}$. Since $I\in J_{\lambda}\subseteq\mathcal{A}_{\lambda}^{\mathrm{prim}}[\prod_{\alpha\in\mathcal{W}(\mathfrak{g}^{\lambda})\smallsetminus\{0\}}\alpha^{-1}]$ by $W^{\lambda}$-invariance of $J_{\lambda}$, $I\in\mathcal{H}_{\lambda}\cap\mathcal{A}_{\lambda}^{\mathrm{prim}}[\prod_{\alpha\in\mathcal{W}(\mathfrak{g}^{\lambda})\smallsetminus\{0\}}\alpha^{-1}]=\mathcal{H}_{\lambda}^{\mathrm{prim}}$. ∎
Lemma 4.13.
The submodule $J_{\lambda}$ is $W_{\lambda}$-invariant.
Proof.
Let $w\in W_{\lambda}$. We let $J_{\lambda}^{w}\coloneqq w\cdot J_{\lambda}$. Then, $J_{\lambda}^{w}$ is $W^{\lambda}$-stable by Lemma 4.3. It is an $\mathcal{A}_{\lambda}^{\mathrm{prim}}$-submodule of $\mathcal{A}_{\lambda}^{\mathrm{prim}}[\prod_{\alpha\in\mathcal{W}(\mathfrak{g})\smallsetminus\{0\}}(w\cdot\alpha)^{-1}]=\mathcal{A}_{\lambda}^{\mathrm{prim}}[\prod_{\alpha\in\mathcal{W}(\mathfrak{g})\smallsetminus\{0\}}\alpha^{-1}]$. Moreover, it contains
$k_{w\cdot\mu,w\cdot\lambda}$
for $\lambda\not\preceq\mu$. Now, we use that by Lemma 2.13, for $w\in W_{\lambda}$, {±$\{\pm k_{\mu,\lambda}\colon\lambda\not\preceq\mu\}=\{\pm k_{w\cdot\mu,w\cdot\lambda}\mid\lambda\not\preceq\mu\}$ since for $w\in W_{\lambda}$, $\overline{w\cdot\lambda}=\overline{\lambda}$. By minimality of $J_{\lambda}$, we deduce that $J_{\lambda}\subseteq J_{\lambda}^{w}$. By symmetry of the argument, we conclude $J_{\lambda}^{w}=J_{\lambda}$. ∎
Corollary 4.14.
$J_{\lambda}^{W^{\lambda}}$ is $W_{\lambda}$-invariant.
Proof.
This follows directly from Lemmas 4.3 and 4.13. ∎
We let $\mathcal{P}_{\lambda}\subseteq\mathcal{H}_{\lambda}^{\mathrm{prim}}$ be a direct sum complement of $J_{\lambda}^{W^{\lambda}}$:
$\mathcal{H}_{\lambda}^{\mathrm{prim}}=J_{\lambda}^{W^{\lambda}}\oplus\mathcal{P}_{\lambda}.$
Lemma 4.15.
For any $\lambda\in X_{*}(T)$, the image of
$\bigoplus_{\lambda\not\preceq\mu}\mathcal{H}_{\mu}\rightarrow\mathcal{H}_{\lambda}$
is $J_{\lambda}^{W^{\lambda}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$.
Proof.
This is a calculation very similar to that in the proof of Lemma 4.13. Namely, if $f\in\mathcal{H}_{\mu}$ and $g\in\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$, $\mathrm{Ind}_{\mu}(fg)=\mathrm{Ind}_{\mu}(f)g$ by $W^{\lambda}$-invariance of $g$ (Lemma 4.5). Moreover, $fg\in\mathcal{H}_{\mu}$. Therefore, a spanning subset of $J_{\lambda}^{W^{\lambda}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$ can be obtained in this way for $\lambda\not\preceq\mu$. Conversely, if $f\in\mathcal{H}_{\mu}\subseteq\mathcal{A}_{\mu}=\mathcal{A}_{\lambda}$ for $\lambda\not\preceq\mu$, we write $f=\sum_{i=1}^{N}f_{i}\otimes g_{i}$ with $f_{i}\in\mathcal{A}_{\lambda}^{\mathrm{prim}}$ and $g_{i}\in\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$ so that
$\mathrm{Ind}_{\mu}(f)=\sum_{i=1}^{N}\mathrm{Ind}_{\mu,\lambda}(f_{i})\otimes g_{i}$
indeed belongs to $J_{\lambda}^{W^{\lambda}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$. ∎
Lemma 4.16.
For any $\lambda\in X_{*}(T)$, the map
$\bigoplus_{\mu\preceq\lambda}\mathcal{P}_{\mu}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\mu})\xrightarrow{\bigoplus_{\mu\prec\lambda}\mathrm{Ind}_{\mu,\lambda}}\mathcal{H}_{\lambda}$
is surjective.
Proof.
We proceed by induction on $\lambda\in X_{*}(T)$. If $\lambda$ is minimal, then $J_{\lambda}=\{0\}$ and so $\mathcal{P}_{\lambda}=\mathrm{H}^{*}(V^{\lambda}/(G^{\lambda}/G_{\lambda}))$. By Proposition 3.7, the map
$\mathcal{P}_{\lambda}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})\rightarrow\mathrm{H}^{*}(V^{\lambda}/G^{\lambda})$
of the lemma is indeed surjective (as it even is an isomorphism).
Now, let $\lambda\in X_{*}(T)$. We assume that the map of the lemma is surjective for any $\lambda^{\prime}\preceq\lambda$. By associativity of the induction (Proposition 2.7), the image of the map of the lemma coincides with the image of the composition
$\left(\bigoplus_{\mu\prec\lambda}\bigoplus_{\mu^{\prime}\prec\mu}\mathcal{P}_{\mu^{\prime}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\mu^{\prime}})\right)\oplus\left(\mathcal{P}_{\lambda}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})\right)\xrightarrow{I\oplus\mathrm{id}}\left(\bigoplus_{\mu\prec\lambda}\mathcal{H}_{\mu}\right)\oplus\left(\mathcal{P}_{\lambda}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})\right)\rightarrow\mathcal{H}_{\lambda}$
where the morphism $I$ is $\bigoplus_{\mu\prec\lambda}\bigoplus_{\mu^{\prime}\prec\mu}\mathrm{Ind}_{\mu^{\prime},\mu}$. By induction hypothesis,
$\left(\bigoplus_{\mu\prec\lambda}\bigoplus_{\mu^{\prime}\prec\mu}\mathcal{P}_{\mu^{\prime}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\mu^{\prime}})\right)\rightarrow\left(\bigoplus_{\mu\prec\lambda}\mathcal{H}_{\mu}\right)$
is surjective and by Lemma 4.15, the image of
$\left(\bigoplus_{\mu\prec\lambda}\mathcal{H}_{\mu}\right)\rightarrow\mathcal{H}_{\lambda}$
is $J_{\lambda}^{W^{\lambda}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$, of which $\mathcal{P}_{\lambda}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$ is a direct sum complement by definition of $\mathcal{P}_{\lambda}$. This proves the surjectivity. ∎
For $y\in\mathfrak{t}$, we let $\mathfrak{g}^{y}=\{\xi\in\mathfrak{g}\mid[y,\xi]=0\}$ and $V^{y}=\{v\in V\mid y\cdot v=0\}$.
Lemma 4.17.
For any $y\in\mathfrak{t}$, there exists $\lambda\in X_{*}(T)$ such that $\mathfrak{g}^{\lambda}=\mathfrak{g}^{y}$ and $V^{\lambda}=V^{y}$.
Proof.
Via the natural embeddings $X_{*}(T)\subseteq\mathfrak{t}_{\mathbf{Q}}$ and $X^{*}(T)\subseteq\mathfrak{t}_{\mathbf{Q}}^{*}$, the natural pairing $X_{*}(T)\times X^{*}(T)\rightarrow\mathbf{Z}$ corresponds to the natural pairing $\mathfrak{t}\times\mathfrak{t}^{*}\rightarrow\mathbf{Q}$. We let
$\mathcal{W}_{1}\coloneqq\{\alpha\in\mathcal{W}^{\prime}(V)\cup\mathcal{W}^{\prime}(\mathfrak{g})\mid\langle y,\alpha\rangle=0\}$
and
$\mathcal{W}_{2}\coloneqq\{\alpha\in\mathcal{W}^{\prime}(V)\cup\mathcal{W}^{\prime}(\mathfrak{g})\mid\langle y,\alpha\rangle\neq 0\}$
The vector space $\cap_{\alpha\in\mathcal{W}_{1}}\alpha^{\perp}$ is non-empty, as it contains $y$. The complement of the finite union of rational hyperplanes $U_{y}\coloneqq\cap_{\alpha\in\mathcal{W}_{1}}\alpha^{\perp}\smallsetminus\cup_{\alpha\in\mathcal{W}_{2}}\alpha^{\perp}$ is therefore also non-empty. We let $y^{\prime}\in U_{y}$. For some $N\geqslant 1$, $Ny^{\prime}\in X_{*}(T)\subseteq\mathfrak{t}$. We let $\lambda\coloneqq Ny^{\prime}$. It satisfies the property of the lemma. ∎
Proposition 4.18.
For any $\lambda\in X_{*}(T)$, $\mathcal{P}_{\lambda}$ is a finite-dimensional vector space.
Proof.
By possibly replacing $V$ by $V^{\lambda}$ and $G$ by $G^{\lambda}$, we may assume that $\lambda=0$. It suffices to prove that $J_{\lambda}^{W^{\lambda}}\subseteq\mathcal{H}_{\lambda}^{\mathrm{prim}}$ has finite codimension. If we replace $V$ by $V\times\mathfrak{g}$, then the corresponding module $J_{\lambda}\subseteq\mathcal{A}^{\mathrm{prim}}_{\lambda}[\prod_{\alpha\in\mathcal{W}(\mathfrak{g})\smallsetminus\{0\}}\alpha^{-1}]$ is smaller than the one for $V$, and therefore so is the corresponding module of invariants $J_{\lambda}^{W^{\lambda}}$, and therefore its codimension is bigger. Moreover, in this case, $J_{\lambda}\subseteq\mathcal{A}_{\lambda}^{\mathrm{prim}}$, i.e. there is no need to localise.
We have an inclusion
$\mathcal{H}_{\lambda}^{\mathrm{prim}}/J_{\lambda}^{W^{\lambda}}\rightarrow\mathcal{A}_{\lambda}^{\mathrm{prim}}/J_{\lambda},$
and it therefore suffices to prove that the ideal $J_{\lambda}\subseteq\mathcal{A}_{\lambda}^{\mathrm{prim}}$ has finite codimension. We see $\mathcal{A}_{\lambda}^{\mathrm{prim}}$ as the ring of regular functions on the affine algebraic variety $\mathfrak{t}/\mathfrak{g}_{\lambda}$. We prove that the closed subscheme defined by $J_{\lambda}$ is supported on $\{0\}$. If the ideal $J_{\lambda}$ vanishes on $y\in(\mathfrak{t}/\mathfrak{g}_{\lambda})_{\overline{\mathbf{Q}}}\smallsetminus\{0\}$ we let $y^{\prime}\in\mathfrak{t}$ be a lift of $y$ and $\nu$ be given by Lemma 4.17. It satisfies $\mathcal{W}^{\nu=0}(V)=\{\alpha\in\mathcal{W}(V)\mid\langle y,\alpha\rangle=0\}\subseteq\mathcal{W}(V)$ and $\mathcal{W}^{\nu=0}(\mathfrak{g})=\{\alpha\in\mathcal{W}(\mathfrak{g})\mid\langle y,\alpha\rangle=0\}\subseteq\mathcal{W}(\mathfrak{g})$. Moreover, at least one of these inequalities is strict. Indeed, we have $y^{\prime}\not\in\mathfrak{g}_{\lambda}$, which means that either $y^{\prime}$ is not in the center of $\mathrm{Lie}(G^{\lambda})$ and so $\nu$ defines a strict Levi of $G^{\lambda}$ or $y^{\prime}$ is not in the Lie algebra of the kernel of the action $G^{\lambda}\rightarrow\mathrm{GL}(V^{\lambda})$ in which case $V^{\nu}\subsetneq V^{\lambda}$. Therefore, $\nu\prec\lambda=0$, i.e $\overline{\nu}$ is nontrivial. By definition, we have $\prod_{\alpha\in\mathcal{W}^{\nu>0}(V^{\lambda})}\alpha\in J_{\lambda}$.
We also have $\prod_{\alpha\in\mathcal{W}^{\nu>0}(V^{\lambda})}\langle y,\alpha\rangle\neq 0$ by choice of $\nu$, for any $\alpha\in\mathcal{W}^{\nu>0}(V^{\lambda})$, $\langle\nu,\alpha\rangle=0$ if and only if $\langle y,\alpha\rangle=0$. This is a contradiction since we found an element $\prod_{\alpha\in\mathcal{W}^{\nu>0}(V^{\lambda})}\alpha\neq 0$ of $J_{\lambda}$ not vanishing on $y$. ∎
Lemma 4.19.
Let $G_{\lambda}^{0}$ be the neutral component of $G_{\lambda}$. Let $k=\dim\mathfrak{g}_{\lambda}$. Then, $G_{\lambda}^{0}\cong(\mathbf{C}^{*})^{k}$ for some $k\geqslant 0$ and the natural pullback map $\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})\rightarrow\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda}^{0})$ is an isomorphism.
Proof.
The group $G_{\lambda}$ is a subgroup of the center of $G^{\lambda}$, and so it is a diagonalisable algebraic group. Therefore, its neutral component is a torus. We have a splitting $G\cong G_{\lambda}^{0}\times\pi_{0}(G_{\lambda})$ [humphreys2012linear, Theorem p.104] and so, by the Künneth formula,
$\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})\cong\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda}^{0})\otimes\mathrm{H}^{*}(\mathrm{pt}/\pi_{0}(G_{\lambda})).$
Since $\pi_{0}(G_{\lambda})$ is a finite group, $\mathrm{H}^{*}(\mathrm{pt}/\pi_{0}(G_{\lambda}))\cong\mathrm{H}^{*}(\mathrm{pt})^{\pi_{0}(G_{\lambda})}\cong\mathbf{Q}$, which concludes. ∎
The rest of this section is devoted to the proof of the injectivity of the cohomological integrality map in Theorem 1.1.
For $\lambda\in X_{*}(T)$, we let $\widetilde{J}_{\lambda}\coloneqq J_{\lambda}\mathcal{A}_{\lambda}$. This is a $\mathcal{A}_{\lambda}$-submodule of $\mathcal{A}_{\lambda}^{\mathrm{loc}}$.
Recall the inclusion $\mathcal{H}_{\lambda}=\mathcal{A}_{\lambda}^{W^{\lambda}}\rightarrow\mathcal{A}_{\lambda}$.
Lemma 4.20.
The elements $\alpha\in\mathcal{W}^{\lambda>0}(V)\cup\mathcal{W}^{\lambda>0}(\mathfrak{g})$ (or more generally, $\alpha\in X^{*}(T)$ such that $\langle\lambda,\alpha\rangle\neq 0$), seen as elements of $\mathcal{A}_{\lambda}$, are not zero divisors in the quotient $\mathcal{A}_{\lambda}$-module $\mathcal{A}_{\lambda}^{\mathrm{loc}}/\widetilde{J}_{\lambda}$.
Proof.
Let $\alpha\in\mathcal{W}^{\lambda>0}(V)\cup\mathcal{W}^{\lambda>0}(\mathfrak{g})$. We see $\alpha\in\mathfrak{t}^{*}$ and $\langle\lambda,\alpha\rangle\neq 0$ implies that $\alpha$ does not vanish on $\mathfrak{g}_{\lambda}$ since it takes a nonzero value on $\left(\frac{\mathrm{d}}{\mathrm{d}t}\lambda(t)\right)_{t=0}\in\mathfrak{g}_{\lambda}$. Therefore, the projection $\alpha^{\prime}\in\mathfrak{g}_{\lambda}^{*}$ of $\alpha$ with respect to the chosen (or any) decomposition $\mathfrak{t}^{*}\cong(\mathfrak{t}/\mathfrak{g}_{\lambda})^{*}\oplus\mathfrak{g}_{\lambda}^{*}$ is nonzero.
We choose a basis $(x_{1},\ldots,x_{\dim\mathfrak{g}_{\lambda}})$ of $\mathfrak{g}_{\lambda}^{*}$ such that $x_{1}=\alpha^{\prime}$ and we order the monomials by the lexicographic order so that $x_{1}\geqslant x_{2}\geqslant\ldots$. Let $g\in\mathcal{A}_{\lambda}^{\mathrm{loc}}\cong\mathcal{A}_{\lambda}^{\mathrm{prim},\mathrm{loc}}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$. It may be written in a unique way (see also Lemma 4.19)
(4.1) $g=\sum_{(\nu_{1},\ldots,\nu_{\dim\mathfrak{g}_{\lambda}})\in\mathbf{N}^{\dim\mathfrak{g}_{\lambda}}}g_{\nu}x^{\nu}$
with $g_{\nu}\in\mathcal{A}_{\lambda}^{\mathrm{prim},\mathrm{loc}}$. We have an equivalence between the following two statements:
  1. (1)
    $g\not\in\widetilde{J}_{\lambda}$,
  2. (2)
    $g_{\nu}\not\in J_{\lambda}$ for some $\nu\in\mathbf{N}^{\dim\mathfrak{g}_{\lambda}}$.
The implication $(1)\Rightarrow(2)$ is immediate as if all $g_{\nu}$ belong to $J_{\lambda}$, then $g$ would be in $\widetilde{J}_{\lambda}$. The reverse implication $(2)\Rightarrow(1)$ is also true by the unicity of the decomposition (4.1) of $g$.
Assume that $g\not\in\widetilde{J}_{\lambda}$. We need to show that for $\alpha\in\mathcal{W}^{\lambda>0}(V)\cup\mathcal{W}^{\lambda>0}(\mathfrak{g})$, $\alpha g\not\in\widetilde{J}_{\lambda}$. We let $\nu=\max\{\nu^{\prime}\colon g_{\nu^{\prime}}\neq 0\}$. We may assume that $g_{\nu^{\prime}}\not\in J_{\lambda}$. We have
$\alpha g=(\alpha-\alpha^{\prime})g+\alpha^{\prime}g$
and $\alpha-\alpha^{\prime}\in\mathrm{Sym}((\mathfrak{t}/\mathfrak{g}_{\lambda})^{*})$. Therefore, the term of highest degree in the variables $(x_{1},\ldots,x_{\dim\mathfrak{g}_{\lambda}})$ in $\alpha g$ is
$g_{\nu}x_{1}x^{\nu}$
and $g_{\nu}\not\in J_{\lambda}$. By the equivalence above, $\alpha g\not\in\widetilde{J}_{\lambda}$. This concludes. ∎
We let $\mathcal{A}_{\lambda}^{\prime}\coloneqq\mathcal{A}_{\lambda}[\prod_{\alpha\in\mathcal{W}^{\lambda>0}(V)\cup\mathcal{W}^{\lambda>0}(\mathfrak{g})}\alpha^{-1}]$, $\mathcal{A}^{\prime\mathrm{loc}}\coloneqq\mathcal{A}_{\lambda}[\alpha^{-1}\colon\alpha\in\mathcal{W}(\mathfrak{g})\smallsetminus\{0\}]$ and $\mathcal{H}_{\lambda}^{\prime}\coloneqq(\mathcal{A}_{\lambda}^{\prime})^{W^{\lambda}}$. We have the localisation maps
$L\colon\mathcal{A}_{\lambda}\rightarrow\mathcal{A}_{\lambda}^{\prime},\quad L\colon\mathcal{H}_{\lambda}\rightarrow\mathcal{H}_{\lambda}^{\prime}.$
We let We let $\widetilde{J}_{\lambda}^{\prime}\coloneqq\mathcal{A}_{\lambda}^{\prime}L(\widetilde{J}_{\lambda})$ be the localised $\mathcal{A}_{\lambda}^{\prime}$-module inside $\mathcal{A}_{\lambda}^{\prime\mathrm{loc}}$.
Lemma 4.21.
The localisation maps
$L\colon\mathcal{A}_{\lambda}^{\mathrm{loc}}/\widetilde{J}_{\lambda}\rightarrow\mathcal{A}_{\lambda}^{\prime\mathrm{loc}}/\widetilde{J}_{\lambda}^{\prime},\quad L\colon\mathcal{H}_{\lambda}/(\widetilde{J}_{\lambda})^{W^{\lambda}}\rightarrow\mathcal{H}_{\lambda}^{\prime}/(\widetilde{J}_{\lambda}^{\prime})^{W^{\lambda}}$
are injective.
Proof.
This follows from Lemma 4.20 as we localise by a set of nonzero divisors of the quotient $\mathcal{A}_{\lambda}^{\mathrm{loc}}/\widetilde{J}_{\lambda}$ in $\mathcal{A}_{\lambda}$. ∎
Lemma 4.22.
Let $\lambda,\mu\in X_{*}(T)$ and $\nu$ as in Lemma 2.6. Then,
$k_{\mu}=k_{\nu,\lambda}\frac{\prod_{\alpha\in\mathcal{W}^{\mu>0}(V)\cap\mathcal{W}^{\lambda\neq 0}(V)}\alpha}{\prod_{\alpha\in\mathcal{W}^{\mu>0}(\mathfrak{g})\cap\mathcal{W}^{\lambda\neq 0}(\mathfrak{g})}\alpha}.$
Proof.
This is immediate from the definitions. ∎
We define
$k^{\prime}_{\mu,\lambda}\coloneqq\frac{\prod_{\alpha\in\mathcal{W}^{\mu>0}(V)\cap\mathcal{W}^{\lambda\neq 0}(V)}\alpha}{\prod_{\alpha\in\mathcal{W}^{\mu>0}(\mathfrak{g})\cap\mathcal{W}^{\lambda\neq 0}(\mathfrak{g})}\alpha}.$
This is an invertible element in $\mathcal{A}_{\lambda}^{\prime}$.
Lemma 4.23.
Let $\lambda,\mu\in X_{*}(T)$ and $w$ in $W$ be such that $w\cdot\mu\neq\lambda$ and $\lambda\not\preceq w\cdot\mu$. Then, $w\cdot k_{\mu}\in\widetilde{J}^{\prime}_{\lambda}$.
Proof.
Of course, the presence of $W$ is superfluous since $wk_{\mu}=k_{w\cdot\mu}$ and it suffices to prove that for any $\lambda,\mu\in X_{*}(T)$ such that $\lambda\not\preceq\mu$, $k_{\mu}\in\widetilde{J}^{\prime}_{\lambda}$.
We let $\nu\in X_{*}(T)$ be such that $V^{\nu}=V^{\lambda}\cap V^{\mu}$, $\mathfrak{g}^{\nu}=\mathfrak{g}^{\lambda}\cap\mathfrak{g}^{\mu}$, $(V^{\lambda})^{\nu\geqslant 0}=(V^{\lambda})^{\mu\geqslant 0}$ and $(\mathfrak{g}^{\lambda})^{\nu\geqslant 0}=(\mathfrak{g}^{\lambda})^{\mu\geqslant 0}$ (Lemma 2.6). By assumption, $\nu\prec\lambda$.
We can write, by Lemma 4.23,
$k_{\mu}=k_{\nu,\lambda}\frac{\prod_{\alpha\in\mathcal{W}^{\mu>0}(V)\cap\mathcal{W}^{\lambda\neq 0}(V)}\alpha}{\prod_{\alpha\in\mathcal{W}^{\mu>0}(\mathfrak{g})\cap\mathcal{W}^{\lambda\neq 0}(\mathfrak{g})}\alpha}.$
This element is in $\widetilde{J}^{\prime}_{\lambda}$ since $k_{\nu,\lambda}$ is in ${J}_{\lambda}$, by Lemma 4.11. ∎
Lemma 4.24.
Let $\lambda\in X_{*}(T)$ be such that for some $w\in W$, $\lambda\preceq w\cdot\lambda$. Then, $\overline{w\cdot\lambda}=\overline{\lambda}$.
Proof.
The lemma follows from the inclusions $V^{\lambda}\subseteq V^{w\cdot\lambda}$ and $\mathfrak{g}^{\lambda}\subseteq\mathfrak{g}^{w\cdot\lambda}$ and the equalities of dimensions $\dim V^{\lambda}=\dim V^{w\cdot\lambda}$ and $\dim\mathfrak{g}^{\lambda}=\dim\mathfrak{g}^{w\cdot\lambda}$. ∎
For $\lambda\in X_{*}(T)$, we have an inclusion
$\mathrm{Res}_{\lambda}\colon\mathcal{H}\cong\mathcal{A}^{W}\cong\mathcal{A}_{\lambda}^{W}\rightarrow\mathcal{H}_{\lambda}\cong\mathcal{A}_{\lambda}^{W^{\lambda}}.$
since $\mathcal{A}\cong\mathcal{A}_{\lambda}$ and $W^{\lambda}\subseteq W$ (Lemma 4.2).
Corollary 4.25.
For any $f_{\mu}\in\mathcal{H}_{\mu}$, $\widetilde{\mu}\in\mathscr{P}_{V}/W$ such that $f_{\mu}=0$ if $\lambda\prec w\cdot\mu$ for some $w\in W$, we have
$\mathrm{Res}_{\lambda}\left(\sum_{\widetilde{\mu}\in\mathscr{P}_{V}/W}\mathrm{Ind}_{\mu}(f_{\mu})\right)\equiv k_{\lambda}\cdot\sum_{w\in W_{\lambda}/W^{\lambda}}\varepsilon_{V,\lambda}(w)(w\cdot f_{\lambda})\pmod{\widetilde{J}^{\prime}_{\lambda}}.$
Proof.
We calculate:
$\mathrm{Res}_{\lambda}\left(\sum_{\widetilde{\mu}\in\mathscr{P}_{V}/W}\mathrm{Ind}_{\mu}(f_{\mu})\right)$ $=\mathrm{Res}_{\lambda}\left(\sum_{\widetilde{\mu}\in\mathscr{P}_{V}/W}\frac{1}{\#W^{\mu}}\sum_{w\in W}(w\cdot f_{\mu})(w\cdot k_{\mu})\right)$
[…]
$=\frac{1}{\#W^{\lambda}}k_{\lambda}\cdot\sum_{w\in W_{\lambda}}\varepsilon_{V,\lambda}(w)(w\cdot f_{\lambda})\quad\text{by definition of the sign $\varepsilon_{V,\lambda}$}.$
Proposition 4.26.
The map
$\bigoplus_{\widetilde{\mu}\in\mathscr{P}_{V}/W}(\mathcal{P}_{\mu}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\mu}))^{\varepsilon_{V,\mu}}\rightarrow\mathrm{H}^{*}(V/G)$
is injective.
Proof.
Let $f_{\mu}\in(\mathcal{P}_{\mu}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\mu}))^{\varepsilon_{V,\mu}}$, $\widetilde{\mu}\in\mathscr{P}_{V}/W$.
We let $T\coloneqq\sum_{\widetilde{\mu}\in\mathscr{P}_{V}/W}\mathrm{Ind}_{\mu}(f_{\mu})$. We assume that not all $f_{\mu}$’s are zero, and we have to prove that $T\neq 0$. We let $\widetilde{\lambda}\in\mathscr{P}_{V}/W$ be such that $f_{\lambda}\neq 0$ and for any $\widetilde{\mu}\in\mathscr{P}_{V}/W$ such that $f_{\mu}\neq 0$, for any $w\in W$, $\lambda\not\preceq\mu$ (otherwise, we may replace $\lambda$ by $\mu$, etc.).
By Corollary 4.25, we have
$\mathrm{Res}_{\lambda}\left(\sum_{\widetilde{\mu}\in\mathscr{P}_{V}/W}\mathrm{Ind}_{\mu}(f_{\mu})\right)\equiv k_{\lambda}\cdot\sum_{w\in W_{\lambda}/W^{\lambda}}\varepsilon(w)(w\cdot f_{\lambda})\equiv k_{\lambda}\#(W_{\lambda}/W^{\lambda})f_{\lambda}\pmod{\widetilde{J}^{\prime}_{\lambda}}.$
and since $f_{\lambda}\not\in\widetilde{J}^{\prime}_{\lambda}$, since $f_{\lambda}\in\mathcal{P}_{\lambda}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\lambda})$, and $k_{\lambda}$ is invertible in $\mathcal{A}_{\lambda}^{\prime}$, $\mathrm{Res}_{\lambda}(T)\neq 0$. Therefore, $T\neq 0$, proving injectivity. ∎
Corollary 4.27.
The map
$\bigoplus_{\widetilde{\mu}}(\mathcal{P}_{\mu}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\mu}))^{\varepsilon_{V,\mu}}\xrightarrow{\bigoplus_{\widetilde{\mu}\in\mathscr{P}_{V}/W}\mathrm{Ind}_{\mu}}\mathrm{H}^{*}(V/G)$
is an isomorphism.
Proof.
The injectivity is Proposition 4.26. The surjectivity comes from Lemmas 4.16 and 2.15 and the formula
$\mathrm{Ind}_{\mu}(f)=\frac{1}{\#W^{\mu}}\sum_{\overline{w}\in W/W_{\mu}}w^{\prime}\cdot\left(\sum_{w\in W_{\mu}}\varepsilon_{V,\mu}(w)w\cdot f\right)(w^{\prime}\cdot k_{\mu})$
deduced from Propositions 2.9 and 4.8 which implies, using Proposition 4.10, that $\mathrm{Ind}_{\mu}(f)$ vanishes if $f$ is not in the $\varepsilon_{V,\mu}$-isotypic component of $\mathcal{P}_{\mu}\otimes\mathrm{H}^{*}(\mathrm{pt}/G_{\mu})$ for the $W_{\mu}$-action. ∎
4.3. The BPS subspaces
In analogy to the case of quivers [efimov2012cohomological, meinhardt2019donaldson], we may call the subspaces $\mathcal{P}_{\lambda}\subseteq\mathrm{H}^{*}(V^{\lambda}/G^{\lambda})$ the BPS subspaces. Although we lack both qualitative and quantitative properties of these spaces, there are cases when we can prove it vanishes.
Proposition 4.28.
Let $\lambda\in X_{*}(T)$. If $\ker(G^{\lambda}\rightarrow\mathrm{GL}(V^{\lambda}))^{0}\not\subseteq Z(G^{\lambda})$, then $\mathcal{P}_{\lambda}=\{0\}$.
Proof.
By possibly replacing $V$ by $V^{\lambda}$ and $G$ by $G^{\lambda}$, we may assume that $\lambda$ is the trivial cocharacter of $T$. If $\ker(G\rightarrow\mathrm{GL}(V))\not\subseteq Z(G)$, then there exists a cocharacter $\mu\colon\mathbf{G}_{\mathrm{m}}\rightarrow\ker(G\rightarrow\mathrm{GL}(V))$ which does not factor through $Z(G)$. By conjugating $\mu$ by some element $g\in G$, we may assume that $\mu\colon\mathbf{G}_{\mathrm{m}}\rightarrow T\cap\ker(G\rightarrow\mathrm{GL}(V))$. We have $\overline{\mu}\in\mathscr{P}_{V}\smallsetminus\{\overline{0}\}$. Then, the induction diagram reads
[…]
and the induction map $\mathrm{Ind}_{\mu}\colon\mathrm{H}^{*}(V/G^{\mu})\rightarrow\mathrm{H}^{*}(V/G)$ is surjective. Hence, $\mathcal{P}_{0}=\{0\}$. ∎
5. Vanishing bounds
In this section, we prove Proposition 1.4.
First, we have an inclusion $\mathcal{P}_{\lambda}^{k}\subseteq\mathcal{H}_{\lambda}^{k}=\mathrm{H}^{k+d_{\lambda}}(V^{\lambda}/G^{\lambda})$, and the latter vector space is zero for $k<-d_{\lambda}=\dim G^{\lambda}-\dim V^{\lambda}$. The proof of the upper bound uses [efimov2012cohomological, Lemma 3.6] and is analogous to the proof of [efimov2012cohomological, Theorem 3.5]. Namely, [efimov2012cohomological, Lemma 3.6] gives $(\mathcal{A}_{\lambda}^{\mathrm{prim}})^{d}\subseteq J_{\lambda}$ for $d\geqslant\dim V^{\lambda}$ (for the unshifted cohomological gradings), where $(\mathcal{A}_{\lambda}^{\mathrm{prim}})^{d}$ denotes degree $d$ polynomials inside $\mathcal{A}_{\lambda}^{\mathrm{prim}}$.
6. Examples
In this section, we give some explicit examples of cohomological integrality isomorphisms for some choices of pairs $(G,V)$ of a reductive group $G$ and a symmetric representation $V$ of $G$. We give each time the explicit formula for the cohomological integrality isomorphisms. In each case, it is possible to verify by hand that they are indeed isomorphisms.
6.1. $(\mathbf{C}^{*})^{2}\curvearrowright\mathbf{C}^{2}\oplus(\mathbf{C}^{2})^{\vee}$
We let $G\coloneqq(\mathbf{C}^{*})^{2}$ act on $V\coloneqq\mathbf{C}^{2}\oplus(\mathbf{C}^{2})^{\vee}\cong\mathbf{C}^{4}$ by
$(t,u)\cdot(a,b,c,d)\coloneqq(ta,ub,t^{-1}c,u^{-1}d).$
We have $\mathscr{P}_{V}=\{\overline{\lambda_{0}},\overline{\lambda_{1}},\overline{\lambda_{2}},\overline{\lambda_{3}}\}$ where
$\begin{matrix}\lambda_{0}&\colon&\mathbf{G}_{\mathrm{m}}&\rightarrow&(\mathbf{C}^{*})^{2}\\ &&v&\mapsto&(1,1)\end{matrix}$
$\begin{matrix}\lambda_{1}&\colon&\mathbf{G}_{\mathrm{m}}&\rightarrow&(\mathbf{C}^{*})^{2}\\ &&v&\mapsto&(v,1)\end{matrix}$
$\begin{matrix}\lambda_{2}&\colon&\mathbf{G}_{\mathrm{m}}&\rightarrow&(\mathbf{C}^{*})^{2}\\ &&v&\mapsto&(1,v)\end{matrix}$
$\begin{matrix}\lambda_{3}&\colon&\mathbf{G}_{\mathrm{m}}&\rightarrow&(\mathbf{C}^{*})^{2}\\ &&v&\mapsto&(v,v)\end{matrix}$
We have $V^{\lambda_{0}}=V$, $V^{\lambda_{1}}=(0\oplus\mathbf{C})\oplus(0\oplus\mathbf{C}^{\vee})$, $V^{\lambda_{2}}=(\mathbf{C}\oplus 0)\oplus(\mathbf{C}^{\vee}\oplus 0)$ and $V^{\lambda_{3}}=\{0\}$. We can describe the induction maps as follows:
$\begin{matrix}\mathrm{Ind}_{\lambda_{1},\lambda_{0}}&\colon&\mathbf{Q}[x_{1},x_{2}]&\rightarrow&\mathbf{Q}[x_{1},x_{2}]\\ &&f&\mapsto&x_{1}f\end{matrix}$
$\begin{matrix}\mathrm{Ind}_{\lambda_{2},\lambda_{0}}&\colon&\mathbf{Q}[x_{1},x_{2}]&\rightarrow&\mathbf{Q}[x_{1},x_{2}]\\ &&f&\mapsto&x_{2}f\end{matrix}$
$\begin{matrix}\mathrm{Ind}_{\lambda_{3},\lambda_{0}}&\colon&\mathbf{Q}[x_{1},x_{2}]&\rightarrow&\mathbf{Q}[x_{1},x_{2}]\\ &&f&\mapsto&x_{1}x_{2}f\end{matrix}$
$\begin{matrix}\mathrm{Ind}_{\lambda_{3},\lambda_{1}}&\colon&\mathbf{Q}[x_{1},x_{2}]&\rightarrow&\mathbf{Q}[x_{1},x_{2}]\\ &&f&\mapsto&x_{2}f\end{matrix}$
$\begin{matrix}\mathrm{Ind}_{\lambda_{3},\lambda_{2}}&\colon&\mathbf{Q}[x_{1},x_{2}]&\rightarrow&\mathbf{Q}[x_{1},x_{2}]\\ &&f&\mapsto&x_{1}f\end{matrix}$
We have
$\mathcal{P}_{\lambda_{0}}=\mathbf{Q}\subseteq\mathbf{Q}[x_{1},x_{2}],\quad\mathcal{P}_{\lambda_{1}}=\mathbf{Q},\quad\mathcal{P}_{\lambda_{2}}=\mathbf{Q},\quad\mathcal{P}_{\lambda_{3}}=\mathbf{Q}$
and the cohomological integrality isomorphism is
$\begin{matrix}\mathbf{Q}\oplus\mathbf{Q}[x_{1}]\oplus\mathbf{Q}[x_{2}]\oplus\mathbf{Q}[x_{1},x_{2}]&\rightarrow&\mathbf{Q}[x_{1},x_{2}]\\ (a_{0},f(x_{1}),g(x_{2}),h(x_{1},x_{2}))&\mapsto&a_{0}+x_{1}f(x_{1})+x_{2}g(x_{2})+x_{1}x_{2}h(x_{1},x_{2}).\end{matrix}$
6.2. $\mathrm{GL}_{2}(\mathbf{C})\curvearrowright\mathbf{C}^{2}\oplus(\mathbf{C}^{2})^{\vee}$
We let $G\coloneqq\mathrm{GL}_{2}(\mathbf{C})$ act on $V\coloneqq\mathbf{C}^{2}\oplus(\mathbf{C}^{2})^{\vee}$ via
$g\cdot(u,v)\coloneqq(gu,(g^{-1})^{t}v),$
where the superscript $t$ indicates the transpose. We let $T\coloneqq(\mathbf{C}^{*})^{2}$ be the standard torus. The Weyl group is $W=\mathfrak{S}_{2}$. We have $\mathscr{P}_{V}/W=\{\widetilde{\lambda_{0}},\widetilde{\lambda_{1}},\widetilde{\lambda_{2}}\}$ where
$\begin{matrix}\lambda_{0}&\colon&\mathbf{G}_{\mathrm{m}}&\rightarrow&(\mathbf{C}^{*})^{2}\\ &&v&\mapsto&(1,1)\end{matrix}$
$\begin{matrix}\lambda_{1}&\colon&\mathbf{G}_{\mathrm{m}}&\rightarrow&(\mathbf{C}^{*})^{2}\\ &&v&\mapsto&(v,1)\end{matrix}$
$\begin{matrix}\lambda_{2}&\colon&\mathbf{G}_{\mathrm{m}}&\rightarrow&(\mathbf{C}^{*})^{2}\\ &&v&\mapsto&(v,v)\end{matrix}$
Therefore, we have $V^{\lambda_{0}}=V$, $V^{\lambda_{1}}=(0\oplus\mathbf{C})\oplus(0\oplus\mathbf{C}^{\vee})$ and $V^{\lambda_{2}}=\{0\}$. We can describe the induction maps as follows:
$\begin{matrix}\mathrm{Ind}_{\lambda_{1},\lambda_{0}}&\colon&\mathbf{Q}[x_{1},x_{2}]&\rightarrow&\mathbf{Q}[x_{1},x_{2}]^{\mathfrak{S}_{2}}=\mathbf{Q}[x_{1}+x_{2},x_{1}x_{2}]\\ &&f&\mapsto&\frac{x_{1}}{x_{1}-x_{2}}f(x_{1},x_{2})+\frac{x_{2}}{x_{2}-x_{1}}f(x_{2},x_{1})\end{matrix}$
$\begin{matrix}\mathrm{Ind}_{\lambda_{2},\lambda_{0}}&\colon&\mathbf{Q}[x_{1},x_{2}]&\rightarrow&\mathbf{Q}[x_{1},x_{2}]^{\mathfrak{S}_{2}}=\mathbf{Q}[x_{1}+x_{2},x_{1}x_{2}]\\ &&f&\mapsto&\frac{x_{1}x_{2}}{x_{1}-x_{2}}f(x_{1},x_{2})+\frac{x_{1}x_{2}}{x_{2}-x_{1}}f(x_{2},x_{1})\end{matrix}$
$\begin{matrix}\mathrm{Ind}_{\lambda_{2},\lambda_{1}}&\colon&\mathbf{Q}[x_{1},x_{2}]&\rightarrow&\mathbf{Q}[x_{1},x_{2}]\\ &&f&\mapsto&x_{2}f\end{matrix}$
The induction $\mathrm{Ind}_{\lambda_{1},\lambda_{0}}$ is surjective, since
$x_{1}+x_{2}=\mathrm{Ind}_{\lambda_{1},\lambda_{0}}(x_{1})$
and
$x_{1}x_{2}=\mathrm{Ind}_{\lambda_{1},\lambda_{0}}\left(\frac{1}{2}x_{1}x_{2}\right).$
A direct sum complement of the image of $\mathrm{Ind}_{\lambda_{2},\lambda_{1}}$ is $\mathbf{Q}[x_{1}]=\mathcal{P}_{\lambda_{1}}\otimes\mathbf{Q}[x_{1}]$.
Therefore, we have $\mathcal{P}_{\lambda_{0}}=0$, $\mathcal{P}_{\lambda_{1}}=\mathbf{Q}$, $\mathcal{P}_{\lambda_{2}}=\mathbf{Q}$ and the cohomological integrality isomorphism reads
$\begin{matrix}\mathbf{Q}[x_{1}]\oplus\mathbf{Q}[x_{1},x_{2}]^{\mathrm{sgn}}&\rightarrow&\mathbf{Q}[x_{1},x_{2}]^{\mathfrak{S}_{2}}\\ (f(x_{1}),g(x_{1},x_{2}))&\mapsto&\frac{x_{1}f(x_{1})-x_{2}f(x_{2})}{x_{1}-x_{2}}+2x_{1}x_{2}\frac{g(x_{1},x_{2})}{x_{1}-x_{2}}\end{matrix}$
where $\mathbf{Q}[x_{1},x_{2}]^{\mathrm{sgn}}$ is the sign-isotypic component, since the kernel $\frac{x_{1}x_{2}}{x_{1}-x_{2}}$ changes sign when we exchange the variables $x_{1},x_{2}$.
6.3. $\mathrm{GL}_{2}(\mathbf{C})\curvearrowright(\mathbf{C}^{2}\oplus(\mathbf{C}^{2})^{\vee})^{g}$
We generalize the previous situation by considering the diagonal action of $\mathrm{GL}_{2}(\mathbf{C})$ on $(\mathbf{C}^{2}\oplus(\mathbf{C}^{2})^{\vee})^{g}$. We can describe the induction maps as follows, analogously to §6.2.
$\begin{matrix}\mathrm{Ind}_{\lambda_{1},\lambda_{0}}&\colon&\mathbf{Q}[x_{1},x_{2}]&\rightarrow&\mathbf{Q}[x_{1},x_{2}]^{\mathfrak{S}_{2}}=\mathbf{Q}[x_{1}+x_{2},x_{1}x_{2}]\\ &&f&\mapsto&\frac{x_{1}^{g}}{x_{1}-x_{2}}f(x_{1},x_{2})+\frac{x_{2}^{g}}{x_{2}-x_{1}}f(x_{2},x_{1})\end{matrix}$
$\begin{matrix}\mathrm{Ind}_{\lambda_{2},\lambda_{0}}&\colon&\mathbf{Q}[x_{1},x_{2}]&\rightarrow&\mathbf{Q}[x_{1},x_{2}]^{\mathfrak{S}_{2}}=\mathbf{Q}[x_{1}+x_{2},x_{1}x_{2}]\\ &&f&\mapsto&\frac{x_{1}^{g}x_{2}^{g}}{x_{1}-x_{2}}f(x_{1},x_{2})+\frac{x_{1}^{g}x_{2}^{g}}{x_{2}-x_{1}}f(x_{2},x_{1})\end{matrix}$
$\begin{matrix}\mathrm{Ind}_{\lambda_{2},\lambda_{1}}&\colon&\mathbf{Q}[x_{1},x_{2}]&\rightarrow&\mathbf{Q}[x_{1},x_{2}]\\ &&f&\mapsto&x_{2}^{g}f\end{matrix}$
We have $\mathcal{P}_{\lambda_{0}}=\bigoplus_{j=0}^{g-2}\mathbf{Q}(x_{1}+x_{2})^{j}$, $\mathcal{P}_{\lambda_{1}}=\bigoplus_{j=0}^{g-1}\mathbf{Q}x_{2}^{j}$, $\mathcal{P}_{\lambda_{2}}=\mathbf{Q}$ and the cohomological integrality isomorphism reads
$\begin{matrix}\mathcal{P}_{\lambda_{0}}\oplus(\mathcal{P}_{\lambda_{1}}\otimes\mathbf{Q}[x_{1}])\oplus\mathbf{Q}[x_{1},x_{2}]^{\mathrm{sgn}}&\rightarrow&\mathbf{Q}[x_{1},x_{2}]^{\mathfrak{S}_{2}}\\ (f(x_{1},x_{2}),g(x_{1},x_{2}),h(x_{1},x_{2}))&\mapsto&f(x_{1},x_{2})+\frac{x_{1}^{g}g(x_{1},x_{2})-x_{2}^{g}g(x_{2},x_{1})}{x_{1}-x_{2}}+2x_{1}^{g}x_{2}^{g}\frac{h(x_{1},x_{2})}{x_{1}-x_{2}}.\end{matrix}$
6.4. $\mathrm{SL}_{2}(\mathbf{C})\curvearrowright\mathbf{C}^{2}$
We let $G\coloneqq\mathrm{SL}_{2}(\mathbf{C})$ act on $\mathbf{C}^{2}$ via the natural action. It is a symmetric representation. We have $W=\mathfrak{S}_{2}$. We let $T\subseteq\mathrm{SL}_{2}(\mathbf{C})$ be the maximal torus of diagonal matrices. We have $\mathscr{P}_{V}/W=\{\widetilde{\lambda_{0}},\widetilde{\lambda_{1}}\}$ where
$\begin{matrix}\lambda_{0}&\colon&\mathbf{G}_{\mathrm{m}}&\rightarrow&T\\ &&v&\mapsto&(1,1)\end{matrix}$
$\begin{matrix}\lambda_{1}&\colon&\mathbf{G}_{\mathrm{m}}&\rightarrow&T\\ &&v&\mapsto&(v,v^{-1}).\end{matrix}$
We have $V^{\lambda_{0}}=V$, $V^{\lambda_{1}}=\{0\}$. We can describe the induction map as follows.
$\begin{matrix}\mathrm{Ind}_{\lambda_{1},\lambda_{0}}&\colon&\mathbf{Q}[x]&\rightarrow&\mathbf{Q}[x^{2}]\\ &&f&\mapsto&\frac{1}{2}(f(x)+f(-x))\,.\end{matrix}$
We have $\mathcal{P}_{\lambda_{1}}=\mathbf{Q}$ and $\mathcal{P}_{\lambda_{0}}=\{0\}$ and the cohomological integrality isomorphism is
$\begin{matrix}\mathbf{Q}[x^{2}]&\rightarrow&\mathbf{Q}[x]\\ f&\mapsto&f\,.\end{matrix}$
6.5. $\mathrm{SL}_{2}(\mathbf{C})\curvearrowright\mathbf{C}^{d}$
Let $V_{d}$ be the $d$-dimensional irreducible representation of $\mathrm{SL}_{2}(\mathbf{C})$. Up to a scaling factor $C_{d}$, the induction morphism is
$\begin{matrix}\mathrm{Ind}_{\lambda_{1},\lambda_{0}}&\colon&\mathbf{Q}[x]&\rightarrow&\mathbf{Q}[x^{2}]\\ &&f&\mapsto&C_{d}x^{\lfloor\frac{d-1}{2}\rfloor}f(x)+C_{d}(-x)^{\lfloor\frac{d-1}{2}\rfloor}f(-x).\end{matrix}$
We have $\mathcal{P}_{\lambda_{1}}=\mathbf{Q}$ and $\mathcal{P}_{\lambda_{0}}=\mathbf{Q}[x^{2}]_{\deg<\lfloor\frac{d-1}{2}\rfloor}$ where $C_{d}\in\mathbf{Q}$ is some constant depending only on $d$.
The cohomological integrality isomorphism is
$\begin{matrix}\mathcal{P}_{\lambda_{0}}\oplus\mathbf{Q}[x]^{\varepsilon_{d}}&\rightarrow&\mathbf{Q}[x]\\ f,g&\mapsto&f+2C_{d}x^{\lfloor\frac{d-1}{2}\rfloor}g\end{matrix}$
where $\mathbf{Q}[x]^{\varepsilon_{d}}=\mathbf{Q}[x^{2}]$ if $\lfloor\frac{d-1}{2}\rfloor$ is even, and $\mathbf{Q}[x]^{\varepsilon_{d}}=x\mathbf{Q}[x^{2}]$ if $\lfloor\frac{d-1}{2}\rfloor$ is odd.
Remark 6.1.
We note, for example in §6.2 and also §6.4, that the integrality morphisms are not isomorphisms when we consider cohomology with integral coefficients instead of rational coefficients. In §6.2, this comes from the presence of the factor $2$ in the cohomological integrality morphism.
6.6. $\mathrm{SL}_{2}(\mathbf{C})\curvearrowright(\mathfrak{sl}_{2}(\mathbf{C}))^{g}$
Let $g\geqslant 0$. We consider the diagonal adjoint action of $\mathrm{SL}_{2}(\mathbf{C})$ on $V\coloneqq(\mathfrak{sl}_{2}(\mathbf{C}))^{g}$. Again, $\mathscr{P}_{V}=\{\overline{\lambda_{1}},\overline{\lambda_{1}}\}$ with $\lambda_{0},\lambda_{1}$ as in §6.4. The induction map reads
$\begin{matrix}\mathrm{Ind}_{\lambda_{1},\lambda_{0}}&\colon&\mathbf{Q}[x]&\rightarrow&\mathbf{Q}[x^{2}]\\ &&f&\mapsto&(2x)^{g-1}f(x)+(-2x)^{g-1}f(-x).\end{matrix}$
We have $V^{\lambda_{1}}\cong\mathbf{C}^{g}$ with the trivial action of $T$. Therefore, $\mathcal{P}_{\lambda_{1}}=\mathbf{Q}$ and $\mathcal{P}_{\lambda_{0}}=\mathbf{Q}[x^{2}]_{\deg<g-1}$. The cohomological integrality isomorphism is
$\begin{matrix}\mathbf{Q}[x]^{\epsilon_{g}}\oplus\mathcal{P}_{0}&\rightarrow&\mathbf{Q}[x^{2}]\ f,g&\mapsto&g(x)+2^{g}x^{g-1}f(x)\end{matrix}$
where $\mathbf{Q}[x]^{\epsilon_{g}}=\mathbf{Q}[x^{2}]$ if $g-1$ is even and $\mathbf{Q}[x]^{\epsilon_{g}}=x\mathbf{Q}[x^{2}]$ is $g-1$ is odd.
Remark 6.2.
This example is an example for which $G_{\lambda_{0}}=\{\pm 1\}$ is a finite group. We see that it does not affect the cohomological integrality isomorphism.
6.7. $G\curvearrowright\mathfrak{g}$
Let $G$ be a reductive group and $\mathfrak{g}$ its Lie algebra with the adjoint action. Then, $\mathscr{P}_{\mathfrak{g}}$ is inn one-to-one correspondence with the set of standard Levi subgroups of $G$ and $\mathscr{P}_{\mathfrak{g}}/W$ is in bijection with conjugacy classes of Levi subgroups of $\mathfrak{g}$. In this case, $\mathbf{Q}[\mathfrak{t}]^{W}$ is a polynomial algebra and for $\lambda\preceq\mu\in X_{*}(T)$, the induction map is
$\begin{matrix}\mathrm{Ind}_{\lambda,\mu}&\colon&\mathbf{Q}[\mathfrak{t}]^{W^{\lambda}}&\rightarrow&\mathbf{Q}[\mathfrak{t}]^{W^{\mu}}\\ &&f&\mapsto&\sum_{w\in W^{\mu}/W^{\lambda}}w\cdot f.\end{matrix}$
Therefore, the induction maps are surjective. This implies that $\mathcal{P}_{\lambda}=0$ except if $\lambda$ is a generic cocharacter (i.e. $\mathfrak{g}^{\lambda}$ is a torus), in which case $\mathcal{P}_{\lambda}=\mathbf{Q}$. The cohomological integrality isomorphism is then
$\begin{matrix}\mathbf{Q}[\mathfrak{t}]^{W}&\rightarrow&\mathbf{Q}[\mathfrak{t}]^{W}\\ f&\mapsto&(\#W)\cdot f\,.\end{matrix}$
6.8. Topology of the algebra of invariants
Computing the polynomial invariants of binary forms is a long-standing problem in invariant theory, solved for binary forms with small degrees, and still challenging in higher degrees.
The sheafified version of our cohomological integrality theorem, which is the subject of a forthcoming work, would give an algorithm to compute the intersection cohomology of the GIT quotients $V_{d}/\!\!/\mathrm{SL}_{2}$ where $V_{d}$ is the $d$-dimensional irreducible representation of $\mathrm{SL}_{2}(\mathbf{C})$.
For $d\in\mathbf{N}$, one can define an $\mathrm{SL}_{2}(\mathbf{C})$-action on the space of binary forms of the form
$b(x,y)=\sum_{j=0}^{d}a_{i}x^{i}y^{d-i}$
by making $\mathrm{SL}_{2}(\mathbf{C})$ act on the coordinates $(x,y)$ via the natural representation. This induces an action of $\mathrm{SL}_{2}(\mathbf{C})$ on the polynomial algebra $\mathbf{C}[a_{i},1\leqslant i\leqslant d]$. One wants to compute the invariant ring $\mathbf{C}[a_{i},1\leqslant i\leqslant d]^{\mathrm{SL}_{2}(\mathbf{C})}$. Let $V_{d}$ be the irreducible $d$-dimensional representation of $\mathrm{SL}_{2}(\mathbf{C})$. In modern terms, we try to compute the geometric invariant theory quotient $V_{d}/\!\!/\mathrm{SL}_{2}(\mathbf{C})$. We may give answers to this problem in small degrees.
In our situation, motivated by integrality results, we also want to determine the stable locus inside $V_{d}$, that is the open subset of closed $\mathrm{SL}_{2}(\mathbf{C})$-orbits with finite stabilizer.
6.8.1. Invariants of linear binary forms
Let $V_{2}=\mathbf{C}^{2}$ be the natural representation of $\mathrm{SL}_{2}(\mathbf{C})$. Then, we calculate easily $V_{2}/\!\!/\mathrm{SL}_{2}(\mathbf{C})=\mathrm{pt}$ since the action of $\mathrm{SL}_{2}(\mathbf{C})$ on $V_{2}$ has two orbits, $\{0\}$ and $V_{2}\smallsetminus\{0\}$, of those only $\{0\}$ is closed. The stabilizer of the closed orbit $\{0\}$ is $\mathrm{SL}_{2}(\mathbf{C})$, which is not finite. Therefore, the stable locus is empty.
6.8.2. Invariants of binary quadratic forms
We let $V_{3}$ be the $3$-dimensional irreducible representation of $\mathrm{SL}_{2}(\mathbf{C})$. There is single polynomial invariant for binary quadratic forms $ax^{2}+bxy+cy^{2}$, which is the discriminant
$D=b^{2}-4ac.$
Therefore, we have $V_{2}/\!\!/\mathrm{SL}_{2}(\mathbf{C})\cong\mathbf{C}$. Since $\dim V_{3}/\mathrm{SL}_{2}(\mathbf{C})=0$ and $\dim\mathbf{C}=1$, the stable locus is empty.
6.8.3. Invariants of binary cubic forms
We let $V_{4}$ be the $4$-dimensional irreducible representation of $\mathrm{SL}_{2}(\mathbf{C})$. There is one polynomial invariant for binary cubic forms $ax^{3}+3bx^{2}y+3cxy^{2}+dy^{3}$, which is the discriminant
$D=3b^{2}c^{2}+6abcd-4b^{3}d-4c^{3}a-a^{2}d^{2}.$
Therefore, $V_{4}/\!\!/\mathrm{SL}_{2}(\mathbf{C})\cong\mathbf{C}$. Moreover, since $\dim V_{4}/\mathrm{SL}_{2}(\mathbf{C})=1=\dim\mathbf{C}$, the stable locus is non-empty. By $\mathbf{C}^{*}$-equivariance, the stable locus is the open locus of binary cubic forms of non-zero discriminant.
We notice that $\dim\mathrm{IH}(\mathbf{C}^{2})=1=\dim\mathcal{P}_{\lambda_{0}}=1$ where $\mathcal{P}_{\lambda_{0}}$ is given in §6.5 for $d=4$.
6.8.4. Invariants of binary quartic forms
Let $V_{5}$ be the irreducible $5$-dimensional representation of $\mathrm{SL}_{2}(\mathbf{C})$. The ring of invariants of binary quartic forms has two algebraically independent generators $i,j$. Therefore, $V_{5}/\!\!/\mathrm{SL}_{2}(\mathbf{C})\cong\mathbf{C}^{2}$. Moreover, $\dim V_{5}/\mathrm{SL}_{2}(\mathbf{C})=2=\dim\mathbf{C}^{2}$ and therefore, the stable locus is non-empty.
6.8.5. Invariants of binary quintic forms
Let $V_{6}$ be the irreducible $6$-dimensional representation of $\mathrm{SL}_{2}(\mathbf{C})$. There are four polynomial invariants $I_{4},I_{8},I_{12},I_{18}$, of respective degrees $4,8,12,18$. The first three $I_{4},I_{8},I_{12}$ are algebraically independent while the square of the last can be expressed in terms of them:
$I_{18}^{2}=-144I_{12}^{3}+(I_{4}^{3}-72I_{4}I_{8})I_{12}^{2}+(24I_{8}^{2}-6I_{4}^{2}I_{8}^{2})I_{12}+9I_{4}I_{8}^{4}.$
This realizes the GIT quotient $V_{6}/\!\!/\mathrm{SL}_{2}(\mathbf{C})$ as a double cover of $\mathbf{C}^{3}$ ramified over the hypersurface $-144I_{12}^{3}+(I_{4}^{3}-72I_{4}I_{8})I_{12}^{2}+(24I_{8}^{2}-6I_{4}^{2}I_{8}^{2})I_{12}+9I_{4}I_{8}^{4}=0$.
6.8.6. Invariant of binary sextic forms
Let $V_{7}$ be the $7$-dimensional irreducible representation of $\mathrm{SL}_{2}(\mathbf{C})$. The ring of polynomial invariants of binary sextic forms has $5$ generators $I_{2},I_{4},I_{6},I_{10},I_{15}$ of respective degrees $2,4,6,10,15$. The generators $I_{2},I_{4},I_{6},I_{10}$ are algebraically independent and $I_{1}^{2}$ can be expressed in terms of $I_{2},I_{4},I_{6},I_{10}$. This realizes $V_{7}/\!\!/\mathrm{SL}_{2}(\mathbf{C})$ has a double cover of $\mathbf{C}^{4}$ ramified over a hypersurface. Moreover, since $\dim V_{7}/\!\!/\mathrm{SL}_{2}(\mathbf{C})=4=\dim\mathbf{C}^{4}$, the stable locus is non-empty.
[…]